### v1

Abstract
Objective: This study proposes a Natural Language Processing pipeline to evaluate prescription appropriateness in Italian referrals, where reasons for prescriptions are recorded only as free text, complicating automated comparisons with guidelines. The pipeline aims to derive, for the first time, a comprehensive summary of the reasons behind these referrals and a quantification of their appropriateness. While demonstrated in a specific case study, the approach is designed to generalize to other types of examinations.
Methods: Leveraging embeddings from a transformer-based model, the proposed approach clusters referral texts, maps clusters to labels, and aligns these labels with existing guidelines. We present a case study on a dataset of 496,971 referrals, consisting of all referrals for venous echocolordopplers of the lower limbs between 2019 and 2021 in the Lombardy Region. A sample of 1,000 referrals was manually annotated to validate the results.
Results: The pipeline exhibited high performance for referrals’ reasons (Prec=92.43%, Rec=83.28%) and excellent results for referrals’ appropriateness (Prec=93.58%, Rec=91.52%) on the annotated subset. Analysis of the entire dataset identified clusters matching guideline-defined reasons - both appropriate and inappropriate - as well as clusters not addressed in the guidelines. Overall, 34.32% of referrals were marked as appropriate, 34.07% inappropriate, 14.37% likely inappropriate, and 17.24% could not be mapped to guidelines.
Conclusions: The proposed pipeline effectively assessed prescription appropriateness across a large dataset, serving as a valuable tool for health authorities. Findings have informed the Lombardy Region’s efforts to strengthen recommendations and reduce the burden of inappropriate referrals.
keywords:
Natural language processing, Referrals, Prescription appropriateness, Public healthcare system, Quality of healthcare
1 Background and significance
In Italy, the National Health System (NHS) is the primary provider of healthcare services, offering universal coverage, with decentralized administration at the regional level [1]. Despite a strong performance in international rankings, the Italian NHS has faced growing challenges in providing timely access to healthcare services in recent years, leading many citizens to turn to private providers [2].
One significant contributor to these delays is the issue of prescription appropriateness, a concern shared by many advanced healthcare systems [3, 4]. Inappropriate prescriptions have a threefold negative impact on the healthcare system: waste of resources, increased waiting times, and reduced effectiveness of the care provided to citizens, who face delays in obtaining accurate diagnoses [5]. In the Italian NHS, referrals from general practitioners (GPs) or specialized physicians are necessary to book examinations. Assessing prescription appropriateness requires understanding the reasons underlying these referrals.
When Italian physicians fill out a referral, they have to use a standardized form, providing the reason for the prescription in a specific free-text field called Clinical Question. The unstructured nature of this field, coupled with the lack of standardized alternatives, makes it challenging to analyze. Addressing this issue requires Natural Language Processing (NLP) techniques.
The number of studies analyzing texts of Italian referrals is very limited. Venturelli et al. [6] used proprietary commercial software (Clinika VAP [7]), for which the NLP models or algorithms are not disclosed. This software provides an appropriate/inappropriate label for each referral, enabling the quantification of inappropriate referrals. However, it does not provide a clustered analysis of the referrals, with a quantification of the different prescription reasons. To the best of our knowledge, the only other study applying NLP on Italian referrals’ texts is [8], which focused on extracting and analyzing follow-up waiting times from clinical questions to investigate timely access to care.
While several studies have evaluated prescription appropriateness for medications within the Italian NHS using structured data or manual text analysis (e.g., [9, 10, 11]), research on the appropriateness of examinations is more limited [12, 13] and has not incorporated natural language processing.
A few studies developed decision-support systems for guideline adherence, both in primary care [14, 15] and hospital settings [16, 17]. These tools assist physicians through closed-form questions, possibly complemented by keyword matching on the text they write. While effective in specific scenarios [18, 19], such tools are not suited for retrospective analyses of their performance or for quantifying the prevalence of appropriate or inappropriate referrals.
Beyond the Italian context, Haddad et al. [20] analyzed Russian referrals’ texts for prescription appropriateness using labelled data; Sagheb et al. [21] designed a rule-based algorithm to verify guidelines adherence for asthma patients from EHR documents; similarly in [22] a rule-based approach was applied to assess guideline adherence for gout patients. Dimitrov et al. [23] conducted a nationwide evaluation of breast cancer guideline adherence in EHRs using proprietary commercial software (Danny platform [24]), which incorporated rules and manual reviews. Villeana et al. [25], analyzed Chilean referrals’ texts using labelled data to binary classify them for correct priority classes.
These studies, however, are often disease-specific or reliant on labelled data or proprietary software, limiting their broader applicability.
2 Objective
This study proposes an NLP-based pipeline to assess the appropriateness of Italian referrals, developing a model designed for unsupervised analysis across different types of medical examinations.
In 2015, the Lombardy Region, the most populated Italian region with approximately 10 million inhabitants - akin to a small state - issued two resolutions (DGRs) focusing on prescription appropriateness in angiology [26] and gastroenterology [27], two fields where the literature reports high rates of inappropriate prescriptions, both in Italy and internationally [28, 29, 30, 31, 32]. These resolutions established guidelines specifying appropriate and inappropriate clinical questions for a variety of examinations.
Building on these efforts, we analyzed examinations covered by these resolutions, contributing to the issuance of a new resolution in December 2023 that reinforced the previous guidelines [33]. This paper focuses on a case study of one of these examinations: the venous echocolordoppler of the lower limbs. In particular, we analyzed a dataset consisting of all referrals prescribed in the Lombardy Region for this exam between 2019 and 2021, for a total of 496,971 referrals, encompassing 229,509 unique clinical questions.
The analysis presented in this study aims to verify the effectiveness of the 2015 resolutions in enforcing prescriptive appropriateness, providing for the first time an overview of the main reasons why these examinations are prescribed, with an approximate quantification of the number of referrals for each disease/symptom and their alignment on existing guidelines, where possible.
Given the absence of labelled data and the need to uncover reasons not addressed in guidelines, we framed this as a clustering problem. In particular, our pipeline clusters clinical questions by exploiting the embeddings produced by a fine-tuned BERT model [34]. These clusters are summarized and, where possible, mapped to guidelines to determine whether they are appropriate, inappropriate, or not covered by the guidelines.
A subset of 1,000 distinct clinical questions was manually annotated to validate the pipeline. Results are reported for this validation subset to verify the correctness of the pipeline output and for the full dataset to analyze prescription appropriateness at the regional level, including stratifications by year (pre-, during, and post-COVID pandemic), geographical area, and type of prescribing physician. Additional results for another type of examination in the gastroenterological domain are included in Appendix D.
The main contributions of this study are the following:
  1. 1.
    The first open-source, fully unsupervised pipeline for assessing prescription appropriateness from Italian referrals, applicable to any type of examination.
  2. 2.
    The first region-wide analysis in Italy of referrals’ reasons and their appropriateness.
3 Material and Methods
3.1 Data
3.1.1 Referrals dataset
The dataset analyzed in this study is extracted from all referrals for venous echocolordoppler of the lower limbs prescribed in the Lombardy Region between 2019 and 2021 (496,971 referrals). Each referral includes the clinical question, the year, the Local Health Authority (LHA) of the prescribing physician and the type of prescribing physician. Physician types are categorized as general practitioners (GPs) or specialized physicians, while LHAs represent the eight healthcare administrative areas within the Lombardy Region. Figure 1 illustrates the distribution of referrals among these groups.
The analysis focuses on the clinical question, a free-text field typically containing one or few sentences. These texts exhibit frequent spelling and typing errors, heavy use of abbreviations, inconsistent punctuation, and non-standard capitalization. The dataset includes 229,509 distinct clinical questions, but only 15% appear more than once. The median length is 71 characters (IQR=39). Only 189 (0.08%) contain a meaningful ICD code, underscoring the necessity of employing NLP for this analysis.
The clinical questions should not contain personal information about patients or physicians. A preliminary check was performed by the Lombardy Region IT provider, masking any suspect personal information before data extraction.
3.1.2 Manual annotations
To validate the identified clusters, a randomly selected subset of 1,000 distinct clinical questions was manually labelled. The labels correspond to categories representing diseases or symptoms that prompted the referral. During an initial review of this subset, ambiguous cases were identified, and an angiologist was consulted to solve doubts and align clusters with prescription guidelines. Annotation guidelines were subsequently refined, and the entire subset was annotated accordingly. Two other reviews of the annotations corrected remaining errors and clarified uncertainties, ensuring the final labelled dataset was error-free. More details on the annotations are provided in Appendix A.
3.1.3 Prescription guidelines
The guidelines used as a reference to assess referral appropriateness are extracted from [26]. In particular, we extracted the appropriate and inappropriate diseases and symptoms for the venous echocolordoppler of the lower limbs. An English translation is reported in Figure 2.
Figure 1: Referrals distributions across years (A), type of prescribing physicians (B) and LHA (C)
Figure 2: Guidelines for referrals for venous echocolordoppler of the lower limbs in the Lombardy Region. Translated from the Appendix 3 of [26].
3.2 Pipeline for prescription appropriateness
In this section, we present our pipeline to evaluate prescription appropriateness. The process can be divided into two main steps:
  1. 1.
    Identify the reason for the referral: extract the central semantic meaning from the clinical question.
  2. 2.
    Verify appropriateness: determine if the identified reason aligns with established guidelines.
The absence of labels for this type of data prevented us from adopting a supervised approach. A rule-based approach was excluded since we aim to build a pipeline that can generalize to any examination, and it would not be feasible to develop tailored rules for each type of examination. Moreover, we are interested in understanding the reasons for these referrals, which might include additional and/or more detailed categories than those mentioned in the guidelines. Because of this, the problem was addressed as a clustering problem.
Figure 3 depicts blocks composing our pipeline, which are detailed below.
Figure 3: Schema of the pipeline for assessing referrals reasons and appropriateness
3.2.1 Pre-processing
Physicians write referrals during patients’ examinations, with limited time available. This leads them to write clinical questions that are far from well-readable and correct texts. The grammar of the sentences is often poor, there are frequent typing errors, and the punctuation is either not used or used incorrectly. The use of uppercase letters does not have any particular meaning, and the use of abbreviations is very frequent. Even in the single type of examination analyzed in this study, multiple abbreviations can indicate the same disease/symptoms. In the manually labelled subset, 15% of clinical questions were marked as having typing errors and 68% as containing abbreviations.
To address these issues, suitable pre-processing steps were applied:
  1. 1.
    lowercasing
  2. 2.
    punctuation removal
  3. 3.
    abbreviation expansion
  4. 4.
    (partial) typos correction
We removed all punctuation and lowercase sentences due to the inconsistent use of these elements.
[…]
Examples of pre-processing effects are reported in Appendix C.
3.2.2 Text embedding
To appropriately cluster the clinical questions, it is necessary to produce a representation capturing the central semantics of these texts without getting confused by their widely present noise. Because of this, vocabulary-based representations, such as TF-IDF [35], cannot be expected to work well for this task. The main alternatives are static and contextual neural network-based embeddings.
Static word embeddings, such as Word2Vec [36], can capture the semantic similarities between words with similar meanings (e.g., vasculopathy - arteriopathy, phlebitis - thrombophlebitis) differently from vocabulary-based representations. Nevertheless, they might have more difficulty distinguishing between relevant and irrelevant tokens in our sentences since they do not consider the context in which tokens appear.
On the contrary, context-based embeddings produced by transformer models consider the surrounding context while embedding each token, and have often shown superior performances in the recent clinical NLP literature [37, 38, 39, 40]. The problem is particularly challenging in our case due to the limited number of embedding models available for the Italian language, especially for the clinical domain.
To this purpose, the same model used in [41] was adopted to cluster sentences extracted from Italian referrals, and we verified that even with clinical questions, its performances are superior to other models. This model consists of a fine-tuned version of the last Italian version of BERT, Umberto [42], on a dataset of publicly available medical documents in Italian and other languages (translated into Italian) [43].
3.2.3 Clustering
HDBscan was adopted to cluster the embeddings since it does not require eliciting in advance the number of clusters and it allows a not complete assignment of items to clusters. HDBscan requires defining the minimum size of the clusters to identify. We verified that the best results are achieved when this parameter is progressively reduced while iterating the clustering on the previously unclustered data. This approach reduced the number of unclustered points, addressing the varying cluster sizes in our dataset.
An additional issue in the clustering procedure was related to the size of our dataset. On a machine with 32 GB of RAM, it was not possible to run the clustering algorithm on the whole dataset. Because of this, we had to divide it into chunks of 30,000 clinical questions and cluster them separately. The subsequent summarization and mapping steps merged the clusters found in different chunks and corresponding to the same diseases/symptoms.
3.2.4 Cluster summarization
Summarizing the identified clusters is necessary to understand the symptoms or diseases corresponding to each cluster. To automatize this step, we extracted keywords from each cluster. The most frequent words in each cluster were initially selected, considering a decreasing frequency threshold until at least two keywords were found. Among these, we excluded those with high frequency in at least half of the clusters. This allowed us to exclude not only stop words but also frequent words for the specific exam (e.g., leg, lower limbs, right, left in our case study).
3.2.5 Mapping clusters to guidelines
After summarization, the keyword representation of the clusters was used to map them to the guidelines. The process involved:
  1. 1.
    Computing embeddings for each clinical question described in the guidelines.
  2. 2.
    Generating an embedding for each cluster based on its keyword summary, enabling direct comparison with the guideline embeddings.
  3. 3.
    Comparing cluster embeddings with guideline embeddings using cosine similarity. Clusters were assigned to the guideline entry with the highest similarity, provided there is one with similarity >0.50absent0.50>0.50> 0.50.
3.3 Experiments and evaluation
3.3.1 Pipeline evaluation
To validate the pipeline, we compared its output with labels from the manually annotated subset. The pipeline was executed on the entire dataset, and results for the manually labelled portion were extracted. This approach ensures fairness, as running the pipeline solely on the labelled subset (1,000 records) would limit its ability to form meaningful clusters due to the restricted data size.
Manual annotations identified 49 distinct labels, with only 13 occurring more than 10 times. For evaluation purposes, these 13 labels were retained for comparison, while the remaining labels were grouped into the other category, which we expect to align with unclustered elements.
The results were evaluated in terms of precision, recall, and f1-score. Metrics were calculated at two levels: at aggregate level to evaluate the overall pipeline performance across all labels, and at per-label level to evaluate performance for individual labels.
Additionally, we weighted clinical questions by their corresponding referral counts, incorporating the dataset’s real-world distribution into the analysis.
A second type of evaluation is performed to evaluate the quality of prescription appropriateness labels. In this case, we grouped the 13 labels into four categories: appropriate, inappropriate, likely inappropriate and undetermined. In particular, we marked as likely inappropriate some clusters for which the indication was not completely clear (see Results section for details on this), and as undetermined those that were not covered by the guidelines.
In both cases, we compared our pipeline with other types of text representations/clustering algorithms. We considered clustering results obtained by embeddings from the original Umberto model, by embeddings from an Italian Word2Vec model [44], by TF-IDF representations and by the LDA algorithm. See Appendix B for details.
3.3.2 Analysis of the entire dataset
The pipeline is applied to the entire dataset, with results that provide a complete quantification of referrals for each identified category, together with their appropriateness. Moreover, we present stratified results by year (pre-, during, and post-COVID pandemic), Local Health Authority (LHA), and prescribing physician type (general practitioners vs specialists).
4 Results
Table 1: Results of clusters mapped to the labels defined by annotations on the manually labelled test set, computed considering clinical questions (CQ) and referrals (REF). P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
P-CQ R-CQ F1-CQ P-REF R-REF F1-REF
Umberto-E3C-Pipeline 86.81 77.32 79.48 92.43 83.28 86.49
Umberto-Base 82.11 73.62 77.21 87.45 77.81 81.89
Word2Vec 76.21 67.55 71.24 81.67 72.35 76.73
TF-IDF 50.11 41.20 42.43 52.54 44.81 45.45
LDA 27.74 34.04 28.48 19.19 27.57 29.90
Table 2: Results of the cluster mapping for each category, on the manually labelled test set, computed for clinical questions (CQ) and referrals (REF). Labels with support <10absent10<10< 10 were grouped with other. P=Precision, R=Recall, F1=F1-Score. P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
Category P-CQ R-CQ F1-CQ # CQ P-REF R-REF F1-REF # REF
Claudicatio 83.33 45.45 58.82 11 90.00 56.25 69.23 16
Pain 83.61 91.07 87.18 56 86.42 93.33 89.74 75
Edema 85.47 96.15 90.50 104 89.51 97.32 93.25 149
Venous insuff. 94.90 74.50 83.47 200 98.06 88.09 92.81 747
Paresthesis 100.00 81.25 89.66 16 100.00 88.00 93.62 25
Saphenectomy 100.00 92.59 96.15 27 100.00 93.94 96.88 33
(Trombo) Phlebitis 98.70 93.83 96.20 81 99.09 93.16 96.04 117
Thrombosis 96.24 94.04 95.13 218 96.99 94.77 95.87 306
Ulcer 95.83 92.00 93.88 25 97.44 92.68 95.00 41
Varices 41.43 89.23 56.59 65 82.47 98.89 89.83 628
Varices surgery 100.00 42.86 60.00 14 100.00 46.67 63.64 15
Vasculopathy 82.14 44.23 57.50 52 95.33 68.92 80.00 148
Other 65.37 67.94 65.68 131 62.56 70.59 66.33 187
4.1 Pipeline evaluation
Table 1 presents the results of the evaluation performed on the manually annotated subset. This evaluation compares the labels assigned to the clustered data with the ones defined by manual annotation. Clusters not mapped to any label or unclustered clinical questions were assigned to the other category.
The pipeline demonstrated superior performance compared to alternative text representations and clustering algorithms, confirming the importance of contextual representations. The metrics were computed for both clinical questions (CQ) and referrals (REF). The referral-level scores were consistently higher due to the influence of clinical questions that occur more frequently, and that seem to be easier to cluster, being less detailed.
Figure 4: Confusion matrix of the clustering assignment to the labels on the manually annotated test set, computed considering the number of clinical questions (A) and the number of referrals (B).
[…]
The confusion matrices shown in Figure 4 allow for a more detailed analysis of misclassification errors. Errors in varices surgery are often due to misclassifications as varices, a strictly related cluster. Regarding varices, its precision is mainly affected by misclassifications from venous insufficiency and other. Looking at the specific clinical questions manually labelled as venous insufficiency and classified as varices, we noticed that some presented both diagnoses, making this misclassification reasonable. For the remaining classes, misclassified samples fall mainly into the other group, affecting its precision.
Table 3: Results of clusters mapped to the appropriate/inappropriate/likely inappropriate/undetermined categories on the manually labelled test set, computed for clinical questions (CQ) and referrals (REF). P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
P-CQ R-CQ F1-CQ P-REF R-REF F1-REF
Umberto-E3C-Pipeline 89.24 88.27 88.64 93.58 91.52 92.51
Umberto-Base 86.02 80.78 83.13 89.15 84.97 86.55
Word2Vec 80.66 74.29 77.01 85.41 77.54 80.88
TF-IDF 73.51 68.45 67.92 79.44 75.70 74.22
LDA 47.61 48.43 47.83 42.64 46.31 41.31
Table 3 shows the results achieved by grouping labels into the appropriate, inappropriate, likely inappropriate and undetermined categories. The results show higher aggregate performance scores compared to Table 1, indicating that most errors occurred within the same appropriateness category rather than across categories. This suggests that distinguishing between appropriate and inappropriate referrals is simpler, while finer distinctions at the label level are more challenging. A detailed list of all identified clusters, together with their mapping, is reported in Appendix B.
Similar results, including validation on a manually annotated subset, are reported in Appendix D for another type of examination.
4.2 Analysis of the entire dataset
Table 4: Summary of referrals’ clusters, mapped to all labels defined in manual annotations. Green indicates appropriate clusters, red inappropriate clusters, yellow clusters more likely to be inappropriate, and light grey those not mentioned in the guidelines.
Cluster Label # REF %
Venous insufficiency 102,339 20,63
Thrombosis 98,821 19,92
Varices 71,269 14,37
Edema 61,822 12,46
(thrombo)phlebitis 49,537 9,99
Pain 21,131 4,26
Vasculopathy 11,244 2,27
Saphenectomy 8,290 1,67
Varices surgery 7,126 1,44
Diabetes 6,583 1,33
Ulcer 6,272 1,26
Paresthesias 4,856 0,98
Claudicatio 3,570 0,72
Dimer 1,601 0,32
Swelling 1,238 0,25
Respiratory insufficiency 1,060 0,21
Heavy legs 705 0,14
Ischaemia 617 0,12
Stent 612 0,12
Dermatitis 474 0,10
Hypercolesterolaemia 416 0,08
Pregnancy 225 0,05
Dialisis 222 0,04
Surgery 198 0,04
Smoker 193 0,04
Dyspnoea 171 0,03
Tingling 170 0,03
Cold legs 147 0,03
Hypoesthesia 80 0,02
Metabolic syndrome 48 0,01
Bleeding 29 0,01
Other 34,917 7,04
Figure 5: Examples of clustered clinical questions. English translations do not completely reflect the abbreviations and typos of the original Italian texts.
Table 4 extends the analysis on the entire dataset, highlighting the appropriateness labels for each cluster. It can be observed that the distribution of the reasons for the referrals is highly unbalanced, with the first five covering more than 75% of referrals. Varices cluster was assigned to likely inappropriate since this diagnosis is inappropriate unless it is associated with surgery, an element that is often difficult to completely exclude, even for a human reader. Summarizing the table, 34.32% of referrals appear to be appropriate, while 34.07% seems to be not appropriate. The remaining 31.61% referrals can be further divided into 14.37% more likely to be not appropriate, 10.20% that are not mentioned on the guidelines, and 7.04% that cannot be clustered or cannot be mapped to the labels defined by manual annotations. Notably, the other group is limited, while there are clusters for many symptoms and diseases not mentioned in the guidelines.
Figure 5 provides some examples, along with their English translation, of clustered clinical questions.
Figure 6: Prescription appropriateness grouped by LHA (A), type of prescribing physicians (B) and year (C)
Figure 6 reports the level of appropriateness stratified by LHA, type of prescribing physicians and year, providing additional insights into the reasons for inappropriate prescriptions.
5 Discussion
In this study, we analyzed texts of Italian referrals for prescription appropriateness, with a case study on venous echocolordoppler of lower limbs. By introducing a novel approach that clusters referral texts and maps these clusters to existing guidelines, we demonstrated how it is possible to extract meaningful and actionable insights into the reasons behind such prescriptions.
Our work has several strengths. The analysis encompassed 496,971 referrals, covering an entire Italian region with 10 million inhabitants for three years. This large-scale dataset enabled a comprehensive evaluation of prescription practices. Validation was performed on a manually annotated subset, demonstrating that contextual embedding models significantly outperformed simpler clustering techniques in capturing the semantic nuances of referral texts. Analyzing the results in detail, we observed very high performances for most diseases/symptoms, with only two groups with lower recall. Additionally, results on a secondary dataset (reported in Appendix D) confirmed the generalizability of our methodology.
The analysis of the whole dataset allowed us to have, for the first time, a clear picture of referral motivations for this examination. This is a relevant result for public health authorities, considering the absence of structured recording of this information in the Italian NHS.
Our mapping of referrals to guidelines revealed frequent inappropriate uses, as well as numerous reasons not explicitly covered by current guidelines. While individually rare, these reasons cumulatively represent a non-negligible share of referrals, suggesting the need to expand the scope of guidelines. The ambiguity in referrals grouped under varices, where physicians often omit whether surgery is indicated, together with generic reasons like pain and vasculopathy, also points to the importance of requesting more specific information from referring physicians.
Stratification revealed significant geographical variability, suggesting that targeted interventions are needed in specific Local Health Authorities. General practitioners exhibited higher rates of inappropriate referrals compared to specialists. This might have multiple reasons, including larger patient base, more frequent encounter with less severe cases, but also a lower knowledge and/or agreement with guidelines [45, 46].
Temporal trends showed a decline in inappropriate referrals during the initial COVID-19 wave in 2020, followed by an increase in 2021 as the healthcare system normalized. This aligns with literature suggesting that during the pandemic’s peak, physicians prioritized urgent and necessary examinations to manage resource constraints effectively [47, 48, 49, 50].
Despite its strengths, this study has some limitations. Manual annotation was performed on a subset of data due to the large time required, and annotations were conducted by a single annotator, preventing inter-annotator agreement analysis. While valid for most cases, the focus on a single reason per referral overlooks referrals with multiple reasons, which could be addressed through fuzzy clustering.
Another limitation was the lack of access to patient history, which prevented a precise assessment of the appropriateness of certain referrals, such as those for varices, where the presence of a surgery often could not be verified.
It is also essential to note that this analysis does not aim to state the appropriateness of single referrals, for which a comprehensive analysis of the clinical status and history of the patient would be necessary. Instead, it aims to provide a comprehensive, automatized and accurate picture at the regional level.
In future research, we aim to expand the validation of our pipeline to other types of examinations, with larger manually annotated datasets, ensuring broader applicability across healthcare services.
6 Conclusions
This study proposed a two-step procedure for assessing the level of appropriateness of Italian referrals. We presented a case study on the venous echocolordoppler of the lower limbs in the Lombardy Region, demonstrating excellent performance on a validation set and extracting relevant insights for public health authorities. The outcomes of this analysis enabled the Lombardy Region to reinforce existing guidelines for physicians, aiming to reduce the number of inappropriate referrals. By addressing one of the key drivers of inefficiency in the healthcare system, this approach has the potential to alleviate significant burdens on the Italian NHS, lowering waiting times and enhancing the overall effectiveness of healthcare services.
[…]
List of abbreviations
  1. 1.
    NHS: National Health Service
  2. 2.
    NLP: Natural Language Processing
  3. 3.
    GP: General Practitioner
  4. 4.
    EHR: Electronic Health Record
  5. 5.
    ATS/LHA: Agenzia di Tutela della Salute / Local Health Authority
  6. 6.
    GDPR: General Data Protection Regulation
  7. 7.
    IQR: Interquartile Range
  8. 8.
    ICD: International Classification of Diseases
  9. 9.
    BERT: Bidirectional Encoder Representation for Transformers
  10. 10.
    DGR: Decreto della Giunta Regionale (Regional Law)
  11. 11.
    REF: Referral
  12. 12.
    CQ: Clinical Question
Declarations
Ethics approval and Consent to participate
The need for ethics approval is deemed unnecessary according to national regulations. In particular, according to the rules from the Italian Medicines Agency (available at: https://www.gazzettaufficiale.it/eli/id/2024/08/20/24A04320/sg), retrospective studies using administrative databases do not require Ethics Committee protocol approval.
The need for consent to participate is deemed unnecessary according to national regulations. In particular, see the rules of the Italian Privacy Authority available at https://www.garanteprivacy.it/web/guest/home/docweb/-/docweb-display/docweb/9124510 and the European Union regulation n. 2016/679 (GDPR) Art. 1 c. 1 (https://gdpr.eu/article-1-subject-matter-and-objectives-overview/), Art. 4 c.1 (https://gdpr.eu/article-4-definitions/) and Recital 26 (https://gdpr-info.eu/recitals/no-26/).
Availability of data and materials
The data that support the findings of this study are available from Regione Lombardia but restrictions apply to the availability of these data, which were used under license for the current study, and so are not publicly available. Data are however available from the authors upon reasonable request and with permission of Regione Lombardia.
Code availability
The code is available by the authors upon request.
Competing interests
The authors declare that they have no competing interests.
Authors’ contributions
VT: investigation, formal analysis, methodology, software, data curation, writing - original draft AB: resources, data curation, conceptualization ME: resources, data curation, conceptualization OL: resources, data curation, conceptualization FI: conceptualization, methodology, formal analysis, supervision, writing - review & editing
Acknowledgments
The present research is part of the activities of “Dipartimento di Eccellenza 2023-2027”. We thank the Welfare Directorate of Regione Lombardia (Agreement between Politecnico di Milano and Regione Lombardia DGR n. XI/5301-27/09/2021) and, in particular, the working group on waiting list monitoring. We thank Dr. Paola Colombo for the consultation regarding the reasons for angiology referrals and Dr. Marco Soncini for gastroenterology referrals.
Appendix A Manual annotations
A validation set of 1000 clinical questions was manually annotated for the main disease/symptom determining the referrals. The existing guidelines from the Lombardy Region [26] were used as a reference, but labels were not forced to be one of the categories mentioned in the guidelines as appropriate or inappropriate. If none of the categories mentioned in the guidelines fit the reason for the referral, a new category representing the main disease or symptom mentioned in the referral was created. The other category was used only when it was impossible to identify a symptom or a disease determining the referral (e.g., “in-depth examination to actually establish whether there is an ongoing pathology given the symptoms present” or “gfwhdfg”).
In case of multiple diseases or symptoms, the following priority order was defined:
  1. 1.
    Reasons mentioned as appropriate in the guidelines
  2. 2.
    Reasons mentioned as inappropriate among the guidelines
  3. 3.
    Reasons not mentioned in the guidelines
If considering these priority levels, there were still multiple diseases/symptoms, the following priority order was defined:
  1. 1.
    If there is one symptom/disease which looks like the leading cause, prefer it
  2. 2.
    If there are multiple symptoms/diseases mentioned in the guidelines, choose the first one to be mentioned in the text
  3. 3.
    If there are multiple symptoms/diseases not mentioned in the guidelines, prefer those for which a category was already created, otherwise the first one to be mentioned in the text
This procedure, after four iterations, including consultation with an angiologist, defined the labels reported in Table A.1.
Table A.1: Labels derived by manual annotations on the validation set of 1000 clinical questions for venous echocolordoppler of the lower limbs. CQ = Clinical Questions, REF = Referrals.
Label # CQ # REF
Thrombosis 218 306
Venous insufficiency 200 747
Edema 104 149
(Thrombo)phlebitis 81 117
Varices 65 628
Pain 56 75
Vasculopaty 52 148
Safenectomy 27 33
Surgery 26 26
Ulcer 25 41
Paresthesis 16 25
Varices surgery 14 15
Other 13 16
Claudicatio 11 16
Swelling 9 10
Transplanctation 8 8
Teleangectasias 7 7
Cardiopathy 6 15
Dermatitis 6 12
Dimer 6 6
Ischaemia 5 5
Diabetes 4 5
Prosthesis 4 4
Cancer 3 5
Heavy Legs 3 4
Hypertension 3 6
Stent 3 2
Arthritis 2 3
Calcification 2 2
Cold legs 2 2
Dyslipidaemia 2 2
Tingling 2 3
Dialysis 1 1
Dyspnoea 1 1
Embolism 1 2
Haemorrhage 1 1
Hypercholesterolaemia 1 3
Hypertrophy 1 1
Hyposthesia 1 1
Infection 1 1
Lymphangitis 1 1
Metabolic syndrome 1 1
Multiple sclerosis 1 1
Neuropathy 1 1
Pregnancy 1 1
Smoker 1 1
Vascular encephalopathy 1 1
Appendix B Clustering
B.1 Additional details on clustering procedures
For TF-IDF representation, we limited the vocabulary to the top 1,000 most frequent tokens. Clustering was performed with HDBscan and K-Means, obtaining the best results with K-Means. The number of clusters was selected with the knee-elbow method analyzing WSS with k from 2 to 50, which led to k=40.
For LDA, the number of topics (clusters) was selected by analyzing coherence with k from 2 to 50, which led to k=7.
For HDBscan, the initial minimum size was set to 500, selected as the minimum size that led to the emergence of at least 5 clusters, and was progressively halved up to 20.
B.2 Additional clustering results
Table LABEL:tab:clusters-summ presents all the clusters identified for venous echocolordoppler of the lower limbs, with their keyword descriptions, the sizes and the labels they were assigned.
Table B.1: Summary of all clusters identified for echocolordoppler of the lower limbs, including their keyword representation, the label assigned them and the number of distinct clinical questions (CQ) and referrals (REF) that they contain.
CLUSTER KEYWORDS [ITA] CLUSTER KEYWORDS [ENG] CLUSTER LABEL # CQ # REF
claudicatio; riferita claudication; referred claudication 1132 1686
claudicatio; intermittens claudicatio; intermittens claudication 653 1179
claudicatio claudicatio claudicatio 402 530
claudicatio; sospetta; claudicazio claudicatio; suspected; claudication claudicatio 60 126
claudicatio; arti; inferiori claudicatio; limbs; lower claudicatio 41 49
dermatite dermatitis dermatitis 421 474
diabete diabetes diabetes 1540 3249
diabete; mellito diabetes; mellitus diabetes 534 1847
diabete; cardiopatia; mellito; diabetes; heart disease; mellitus diabetes 755 1254
diabetico; diabetica diabetes; diabetic diabetes 209 165
dislipidemia; diabete dyslipidaemia; diabetes diabetes 36 68
dialisi dialysis dialysis 110 222
dimero dimer dimer 1343 1601
dispnea dyspnoea dyspnoea 104 171
dolore pain pain 6343 12051
deambulazione; dolore; difficolta walking; pain; difficulty pain 1294 3884
algia; artralgia; piedi algia; arthralgia; foot pain 1028 2491
polpaccio; dolore calf; pain pain 707 1753
algie; arti; gambe; polpacci; bilaterali; natura; determinare pain; limbs; calves; bilateral; nature; determine pain 160 631
dolore; natura; determinare; algie pain; nature; determine; pain pain 53 126
dolore; popliteo pain; popliteal pain 35 117
algia; covid; gamba pain; covid; leg pain 30 63
dolore; polpaccio; natura; determinare pain; calf; nature; determine pain 12 10
crampi; notturni cramps; nocturnal pain 6 5
edema edema edema 11686 20723
edemi; declivi edema; declivities edema 9491 16473
edema; arti; edemi edema; limbs; edemas edema 3712 7867
arti; inferiori; edemi limbs; lower; edemas edema 528 2486
edema; gamba; piede; polpaccio edema; leg; foot; calf edema 1046 2467
edema; arti; inferiori edema; limbs; lower edema 899 2350
edema; perimalleolare edema; perimalleolar edema 1039 2214
edema; natura; determinare edema; nature; determine edema 1027 2199
edemi; perimalleolari edema; perimalleolar edema 1190 2151
edema; malleolare edema; malleolar edema 935 1136
edemi; vasculopatia; declivi; edema edema; vasculopathy; declivities edema 529 602
edema; flebite edema; phlebitis edema 217 297
arti; bilaterale; piede; edema; succulenza; perimalleolare limbs; bilateral; foot; edema; succulence; perimalleolar edema 117 192
arti; inferiori; edema limbs; lower; edema edema 80 168
arti; inferiori; natura; determinare; edema; edemi limbs; lower; nature; determine; edema edema 80 168
edema; inferiori; arti edema; lower; limbs edema 74 163
edema; sinistra edema; left edema 61 108
claudicatio; edemi; edema; cardiopatia; arti; inferiore; lieve claudication; edema; edema; cardiopathy; limbs; lower; mild edema 32 34
linfedema; arti lymphedema; limbs edema 23 24
emorragia haemorrhage haemorrhage 18 29
formicolio tingling tingling 43 170
freddo cold cold 49 147
fumatore smoker smoker 121 193
gravidanza pregnancy pregnancy 92 225
insufficienza respiratoria respiratory insufficiency respiratory insufficiency 724 1060
venosa; insufficienza venous; insufficiency venous insufficiency 30324 71150
venosa; arti; inferiori; insufficienza; sospetta venous; limbs; lower; insufficiency; suspected venous insufficiency 19532 28831
superficiale; venosa; insufficienza superficial; venous; insufficiency venous insufficiency 276 613
insufficienza; venosa; cronica venous; insufficiency; chronic venous insufficiency 415 506
venosa; safenectomia; insufficienza venous; saphenectomy; insufficiency venous insufficiency 280 311
venosa; claudicatio; insufficienza venous; claudication; insufficiency venous insufficiency 109 180
venosa; arti; insufficienza; varici venous; limb; insufficiency; varicose venous insufficiency 110 112
insufficienza; venosa; arti; inferiori venous; insufficiency; limbs; lower venous insufficiency 23 102
venoso; circolo; deficit; arti; venous; circulation; insufficiency; limbs venous insufficiency 74 102
venosa; insufficienza; arti; inferiori; edemi venous; insufficiency; limbs; lower; edemas venous insufficiency 47 58
insufficienza; venosa venous; insufficiency venous insufficiency 0 0
arti; venosa; inferiori limbs; venous; lower venous insufficiency 0 0
venosa; edemi; declivi; insufficienza venous; edemas; declivity; insufficiency venous insufficiency 0 0
venosa; insufficienza; edemi; arti; inferiori venous; insufficiency; edemas; limbs; lower venous insufficiency 0 0
venosa; cronica venous; chronic venous insufficiency 0 0
insufficienza; venosa; cronica; sospetta venous; insufficiency; chronic; suspected venous insufficiency 264 374
intervento; ginocchio; artroprotesi; valutazione surgery; knee; arthroplasty; evaluation intervention 67 77
lista; trapianto; rene; inserimento list; transplant; kidney; insertion intervention 61 69
lista; trapianto; valutazione; inserimento; rene; sinistra; dialisi; intervento list; transplant; evaluation; insertion; kidney; left; dialysis; intervention intervention 49 52
ipercolesterolemia hypercholesterolaemia hypercholesterolaemia 363 416
ipoestesia hypoesthesia hypoesthesia 31 80
ischemia ischaemia ischaemia 517 617
parestesie paresthesias paresthesias 1834 3965
parestesie; gamba; sinistra paresthesias; leg; left paresthesias 541 891
pesantezza heaviness heaviness 215 475
arti; inferiori; linfedema; pesantezza limbs; lower; lymphedema; heaviness heaviness 25 124
arti; inferiori; pesantezza limbs; lower; heaviness heaviness 18 106
safenectomia saphenectomy saphenectomy 1988 4177
safenectomia; esiti saphenectomy; outcomes saphenectomy 1374 3617
safenectomia; pregressa saphenectomy; previous saphenectomy 342 379
safenectomia; edemi saphenectomy; edemas saphenectomy 41 117
sindrome metabolica metabolic syndrome metabolic syndrome 12 48
stent stent stent 301 612
tromboflebite thrombophlebitis thrombophlebitis 7073 17415
flebite; sospetta phlebitis; suspected thrombophlebitis 5408 12379
flebite; gamba; superficiale phlebitis; leg; superficial thrombophlebitis 3037 8850
varicoflebite; sospetta; tromboflebite varicophlebitis; suspected; thrombophlebitis thrombophlebitis 4546 7739
tromboflebite; superficiale thrombophlebitis; superficial thrombophlebitis 950 1648
gamba; tromboflebite; destra leg; thrombophlebitis; right thrombophlebitis 397 486
tromboflebite; safena thrombophlebitis; saphenous vein thrombophlebitis 386 384
flebite; coscia; superficiale phlebitis; thigh; superficial thrombophlebitis 109 194
flebopatia; sospetta; arti; inferiori; sospetta phlebopathy; suspected; lower; limbs thrombophlebitis 209 175
tromboflebite; poplitea; arti; inferiori thrombophlebitis; popliteal; lower; limbs thrombophlebitis 54 102
flebite; superficiale; gamba phlebitis; superficial; leg thrombophlebitis 48 97
tromboflebite; sospetta; thrombophlebitis; suspected thrombophlebitis 30 68
trombosi thrombosis thrombosis 17060 43558
trombosi; venosa; profonda thrombosis; venous; deep thrombosis 17453 38230
trombosi; vascolare; thrombosis; vascular thrombosis 5086 10180
trombosi; venosa venous; thrombosis thrombosis 1325 3337
venosa; trombosi; profonda; sospetto; segni; clinici; clinico venous; thrombosis; deep; suspected; signs; clinical thrombosis 566 1152
poplitea; trombosi popliteal; thrombosis thrombosis 607 642
venosa; trombosi; poplitea venous; thrombosis; popliteal thrombosis 407 525
trombosi; safena; insufficienza thrombosis; saphenous; insufficiency thrombosis 362 472
segni; clinici; sospetto; clinico; trombosi; venosa; profonda; edema signs; clinical; suspected; clinical; venous; thrombosis; deep; edema thrombosis 342 389
safena; trombosi saphenous vein; thrombosis thrombosis 301 263
post; trombotica post; thrombotic thrombosis 36 73
tumefazione swelling swelling 533 741
tumefazione; gamba swelling; leg swelling 172 448
destro; polpaccio; tumefazione; paziente right; calf; swelling; patient swelling 35 49
ulcera; vascolare ulcer; vascular ulcer 828 3861
ulcera ulcer ulcer 1258 2411
varici varices varices 13999 35939
arti; inferiori; varici limbs; lower; varices varices 7018 18157
venose; varici venous; varices varices 4957 13554
varicosita; stasi varicosities; stasis varices 957 2701
varici; arti; inferiori varices; limbs; lower limbs varices 219 277
vene; varicose; algie veins; varicose; algie varices 121 194
safena; varici saphenous veins; varicose varices 47 136
varicosita; varicosita’ varicosities; varicositis varices 86 97
vascolari; esiti; polmonite; varici; chirurgici; linfangite vascular; outcomes; pneumonia; varices; surgical; lymphangitis varices 60 88
varici; sospette varicose veins; suspected varices 34 63
varicosita; gamba varicosities; leg varices 80 63
varici; intervento varicosities; surgery varices intervention 3144 7126
vasculopatia; periferica vasculopathy; peripheral vasculopathy 3820 8200
vasculopatia vasculopathy; peripheral vasculopathy 769 1886
vasculopatia; diabetico vasculopathy; diabetic vasculopathy 393 408
arteriopatia; obliterante arteriopathy; obliterative vasculopathy 288 272
arteriopatia; obliterante; periferica arteriopathy; obliterative; peripheral vasculopathy 47 117
iperteso; vasculopatico; vasculopatia; dislipidemico hypertensive; vasculopathy; dyslipidemic vasculopathy 17 97
aneurisma; aorta; esiti; popliteo; addominale; aortico aneurysm; aorta; popliteal; abdominal; aortic vasculopathy 42 78
vasculopatia; arti; inferiori vasculopathy; limbs; lower vasculopathy 36 78
vasculopatia; arti; inferiori; polidistrettuale vasculopathy; limb; lower; polydistrict vasculopathy 35 49
vasculopatia; venosa vasculopathy; venous vasculopathy 25 34
arteriopatia arteriopathy vasculopathy 20 25
natura; determinare; periferica nature; determine; peripheral other 3978 10039
natura; determinare; arti; inferiori nature; determine; limbs; lower other 5018 9601
paziente; polsi; periferici patient; wrists; peripheral other 2018 8015
arti; inferiori limbs; lower other 2190 5165
safena; incontinenza saphenous; incontinence other 288 782
ipertrofia hypertrophy other 506 591
tromboembolia; polmonare; esiti thromboembolism; pulmonary; outcomes other 386 379
embolia; polmonare embolism; pulmonary other 110 174
femoro; femorale; paziente; popliteo; sospetta; asse; esiti; venosa femoral; femoral; patient; popliteal; suspected; axis; outcomes; venous other 49 49
stenosi; venoso stenosis; venous other 30 49
dislipidemia; ipertensione dyslipidaemia; hypertension other 25 44
aorta; sottorenale aorta; subrenal other 25 29
Table B.1: *
Appendix C Pre-processing results
Table C.1 reports the list of the 50 most frequent abbreviations that were expanded by our pre-processing, while Table C.2 contains an example list of 50 misspelt words that were detected and corrected during our pre-processing.
Table C.1: Most frequent abbreviations in clinical questions for echocolordoppler of the venous limbs, along with their expansions in Italian and English.
Abbr Freq Expansion Expansion [ENG]
dx 41942 destra right
sx 35400 sinistra left
tvp 35392 trombosi venosa profonda deep vein thrombosis
inf 26644 inferiore inferior
sin 8490 sinistra left
pz 6100 paziente patient
ndd 4555 di natura da determinare of a nature to be determined
$a$ 4079 arti limbs
tvs 3650 trombosi venosa superficiale superficial venous thrombosis
sn 3055 sinistra left
x 2210 per for
k 2170 cancro cancer
ivc 2124 insufficienza venosa cronica chronic venous insufficiency
$i$ 1768 inferiori inferior
tep 1695 tromboembolia polmonare pulmonary thromboembolism
nao 1291 nuovi anticoagulanti orali new oral anticoagulants
ca 1262 cancro cancer
ds 1001 destra right
paz 959 paziente patient
tao 924 terapia anticoagulante orale oral anticoagulant therapy
vgs 911 vena grande safena great saphenous vein
tv 875 trombosi venosa venous thrombosis
tp 852 trombosi profonda deep thrombosis
s 819 sindrome syndrome
dm 800 diabete mellito diabetes mellitus
bil 792 bilaterale bilateral
irc 650 insufficienza renale cronica chronic renal failure
ins 620 insufficienza failure
sup 608 superiore upper
fa 592 fibrillazione atriale atrial fibrillation
v 576 venosa venous
iv 497 insufficienza venosa venous insufficiency
spt 469 sindrome post trombotica post-thrombotic syndrome
npl 451 neoplasia neoplasm
iii 444 inferiori inferior
ep 421 episodio episode
tx 364 trapianto transplantation
ctr 351 controllo control
art 327 tromboembolismo venoso venous thromboembolism
tpv 326 trombosi profonda venosa deep vein thrombosis
vs 307 vena safena saphenous vein
cr 297 cronica chronic
$g$ 269 giorni days
tev 256 tromboembolia venosa venous thromboembolism
pta 251 angioplastica percutanea transluminale percutaneous transluminal angioplasty
ven 235 venosa venous
int 227 intervento surgery
mvc 225 malattia venosa cronica chronic venous disease
ev 220 eventuale possible
$t$ 203 tibio-tarsica tibio-tarsal
Table C.2: Sample of 50 misspelt words with the corresponding replacements applied by our pre-processing.
Original Word Corrected Word Corrected Word [ENG]
iprtensioone ipertensione hypertension
stwnosi stenosi stenosis
areomasica ateromasia atheromasia
aduta caduta fall
clacifiche calcifiche calcifications
lipitimiche lipotimie lipothymias
esignificative significative significant
veriginosa vertiginosa vertiginous
lipotimoco lipotimici lipotimics
emicrani emicrania migraine
cardiopatiia cardiopatia heart disease
30anni anni years
sinndrome sindrome syndrome
vertigicnosa vertiginosa vertiginous
eriosa arteriosa arterial
srteriopatia arteriopatia arteriopathy
carotgidee carotidea carotid
nvertiginosa vertiginosa vertiginous
nlipotimia lipotimici lipothymic
pregressotia pregresso previous
priomo primo first
dislipemua dislipidemia dyslipidaemia
caroltidea carotidea carotid
iportesa ipertesa hypertensive
cerebricolare cerebrovascolare cerebrovascular
neurologoin neurologo neurologist
cronarico coronarica coronary
arteiopatia arteriopatia arteriopathy
bilaterati bilaterale bilateral
aritimica aritmica arrhythmic
bpcoe bpco bpco
extrasisoliche extrasistolia extrasystole
paiente paziente patient
iprtnsione ipertensione hypertension
teriopati arteriopatia arteriopathy
caprtidea carotidea carotid
soszpetta sospetta suspected
dislidemica dislipidemia dyslipidaemia
lipotimicio lipotimici lipothymics
epeisodi episodi episodes
ipertrigliceridiemia ipertrigliceridemia hypertriglyceridaemia
ertiginosa vertiginosa vertigo
iperpensione ipertensione hypertension
neopla neoplasia neoplasia
oderata moderata moderate
cardiopataia cardiopatia cardiopathy
polidistrettuae polidistrettuale polydistrict
dislipidemiaecocolo dislipidemico dyslipidaemic
ipwrtensone ipertensione hypertension
acugene acufene tinnitus
Appendix D Additional results on gastroenterology exam
Additional results related to the flexible endoscope colonoscopy with possible biopsies examination are reported in this section. The dataset for this exam consisted of all referrals filled in in the Lombardy Region between 2019 and 2021, for a total of 407,949 referrals. A random sample of 1,000 distinct clinical questions (corresponding to 3,556 referrals) was annotated in the same way as for the case study presented in the main manuscript and used to validate the results.
Considering 2015 guidelines from Regione Lombardia for this type of exam [27], we marked colon cancer familiarity/prevention and chronic inflammatory bowel disease as likely appropriate. Guidelines state that, in these cases, physicians should carefully consider the patient history and condition to establish if and when the exam should be performed.
Table D.1 reports the detailed results per category, while Table D.2 the summary metrics considering the categories and the appropriate/inappropriate levels. Table D.3 presents the final results on the full dataset, which are stratified by year, type of prescribing physician and LHU in Figure D.1.
Table D.4 details the labels from manual annotation on the test set.
Table D.1: Detailed results of the cluster mapping for each category, on the manually labelled test set for flexible endoscope colonscopy, computed for clinical questions (CQ) and referrals (REF). P=Precision, R=Recall, F1=F1-Score.
Category P-CQ [%] R-CQ [%] F1-CQ [%] # CQ P-REF [%] R-REF [%] F1-REF[%] # REF
Bleeding 100.00 77.91 87.59 163 100.00 97.31 98.64 1748
Polyps 99.35 97.47 98.40 158 99.71 98.55 99.12 344
Family history / prevention 97.56 86.33 91.60 139 98.57 91.39 94.85 302
Colon cancer 96.84 85.98 91.09 107 98.24 89.30 93.56 187
Abdominal pain 88.57 89.42 89.00 104 93.66 89.72 91.65 214
Anaemia 92.75 98.46 95.52 65 95.73 97.39 96.55 115
Chronic inflammatory bowel disease 100.00 89.29 94.34 56 100.00 88.37 93.83 86
Constipation 78.38 80.56 79.45 36 91.61 94.24 92.91 139
Diverticulosis 90.00 81.82 85.71 33 86.67 81.25 83.87 48
Altered bowel 84.85 93.33 88.89 30 92.54 95.38 93.94 65
Chronic diarrhoea 100.00 82.14 90.20 28 100.00 93.46 96.62 107
Diverticulitis 100.00 92.86 96.30 14 100.00 89.29 94.34 28
Other 47.14 98.51 63.77 67 59.93 99.42 74.78 173
Table D.2: Summary metrics on the manually labelled test set for flexible endoscope colonscopy, computed for clinical questions (CQ) and referrals (REF), considering annotations categories or only appropriate/inappropriate/likely appropriate/undetermined. P=Precision, R=Recall, F1=F1-Score.
P-CQ [%] R-CQ [%] F1-CQ [%] P-REF [%] R-REF [%] F1-REF [%]
Annotation categories 90.42 88.78 88.60 93.59 92.70 92.67
Appropriateness levels 92.05 90.96 90.88 94.66 93.96 94.05
Table D.3: Summary of referrals’ clusters for flexible endoscope colonscopy, mapped to all labels defined in manual annotations. Green indicates appropriate clusters, red inappropriate clusters, yellow clusters for which the information might not be sufficient to define the appropriateness but is more likely appropriate, and light grey those not mentioned in the guidelines.
Cluster Label # REF %
polyps 81,111 19.88
occult blood 76,487 18.75
colon cancer familiarity / prevention 43,615 10.69
abdominalgia 31,316 7.68
colon cancer 30,434 7.46
anaemia 23,133 5.67
chronic inflammatory bowel disease 21,470 5.26
constipation 15,109 3.70
diverticulosis 15,058 3.69
chronic diarrhoea 13,533 3.32
altered alvo 10,933 2.68
haemorrhoidal disease 5,379 1.32
diverticulitis 4,669 1.14
suspected colic lesion detected by other investigation technique 3,671 0.90
irritable bowel syndrome 2,023 0.50
surgery 797 0.20
lynch syndrome 787 0.19
other 28,424 6.98
Figure D.1: Prescription appropriateness for flexible endoscope colonscopy grouped by LHA (A), type of prescribing physicians (B) and year (C)
Table D.4: Labels derived by manual annotations on the validation set of 1000 clinical questions for flexible endoscope colonoscopy.
Label # CQ # REF
polyps 158 344
bleeding 122 954
colon cancer 107 187
abdominalgia 104 214
colon cancer familiarity 82 141
anaemia 65 115
prevention 57 161
occult blood 36 789
constipation 36 139
diverticulosis 33 48
altered bowel 30 65
colitis 28 40
chronic diarrhoea 28 107
diverticulitis 14 28
crohn’s disease 12 16
haemorrhoidal disease 9 30
chronic inflammatory bowel disease 8 15
ulcerative rectocolitis 8 15
surgery 5 6
intestinal obstruction 5 10
proctorrhagia 5 5
dyspepsia 4 5
weight loss 3 8
irritable bowel syndrome 3 6
lymphoma 3 4
transplant 3 15
other 2 2
colonpathy 2 7
enteritis 2 3
proctitis 2 4
rectal prolapse 2 12
lynch syndrome 2 2
angiodysplasia 1 1
appendicitis 1 1
celiac disease 1 1
intestinal loops distension 1 1
hepatopathy 1 1
protruding formation 1 1
gastritis 1 2
ileitis 1 1
hypercholesterolaemia 1 1
splenic flexure thickening 1 3
rectal wall thickening 1 2
sigma wall thickening 1 1
behcet’s disease 1 1
liver disease 1 1
pancreatitis 1 1
proctalgia 1 3
sigmoiditis 1 1
suspected lesion 1 33
splenic fissure stenosis 1 1
rectal tenesmus 1 2