### v2

Abstract
Background: Assessing the appropriateness of diagnostic referrals is critical for improving healthcare efficiency and reducing unnecessary procedures. However, this task becomes challenging when referral reasons are recorded only as free text rather than structured codes. In the Italian National Health Service (NHS), referral motivations are systematically entered as unstructured text, making large-scale evaluation against official guidelines difficult to automate and scale. To address this gap, we propose a fully unsupervised Natural Language Processing (NLP) pipeline capable of extracting and evaluating referral reasons without relying on labelled datasets, which are costly and time-consuming to produce.
Methods: Our pipeline leverages Transformer-based embeddings pre-trained on Italian medical texts to semantically cluster referral reasons and assess their alignment with appropriateness guidelines. It operates in an unsupervised setting and is designed to generalize across different examination types. We analyzed two complete regional datasets from the Lombardy Region (Italy), covering all referrals between 2019 and 2021 for venous echocolordoppler of the lower limbs (ECD; n=496,971) and flexible endoscope colonoscopy (FEC; n=407,949). The pipeline was developed on the ECD dataset and tested on the FEC dataset. For both, a random sample of 1,000 referrals was manually annotated to measure performance. Results were also stratified by year, region, and physician type.
Results: The pipeline achieved high performance in identifying referral reasons (Precision=92.43% (ECD), 93.59% (FEC); Recall=83.28% (ECD), 92.70% (FEC)) and appropriateness (Precision=93.58% (ECD), 94.66% (FEC); Recall=91.52% (ECD), 93.96% (FEC)). Inappropriate referrals were significantly more prevalent for ECD (34.1%) than for FEC (5.3%). The analysis revealed geographical, temporal, and physician type variations, including a drop in inappropriate referrals during the COVID-19 pandemic in 2020. Additionally, several referral motivations not covered by current guidelines were identified, highlighting areas for policy improvement. These findings contributed to updates to the regional guidelines in Lombardy.
Conclusions: This study presents a robust, scalable, unsupervised NLP pipeline for assessing referral appropriateness from unstructured texts in large, real-world datasets. It demonstrates how such data can be effectively leveraged, providing public health authorities with a deployable AI tool to monitor practices and support evidence-based policy.
keywords:
Natural language processing, Artificial Intelligence, Machine Learning, Referrals, Unsupervised learning, Examination appropriateness, Guideline adherence, Public healthcare system, Quality of healthcare
pacs:
MSC Classification]68T50
1 Background
In Italy, the National Health Service (NHS) is the primary provider of healthcare services, offering universal coverage, with administration decentralized at the regional level [1]. Despite its strong performance in international rankings, the Italian NHS has faced growing challenges in providing timely access to healthcare services in recent years, with increasing numbers of citizens turning to private providers [2].
One significant contributor to these delays is the issue of referrals appropriateness, a concern shared by many advanced healthcare systems [3, 4]. Inappropriate referrals lead to resource waste, increased waiting times, and delays in accurate diagnoses, ultimately reducing the effectiveness of the healthcare system [5]. Within the Italian NHS, referrals from general practitioners (GPs) and specialists are required to book diagnostic examinations, making it crucial to assess the reasons behind these referrals to ensure their appropriateness.
When issuing a referral, Italian physicians must complete a standardized form, including a free-text field known as the Clinical Question, where they specify the reason for the prescription. The unstructured nature of this field, coupled with the absence of standardized alternatives, poses a major challenge for systematic analysis. Addressing this issue requires Natural Language Processing (NLP) techniques. In this study, we frame the task as a clustering problem: clinical questions are grouped using Transformer-based embeddings [6] to capture semantic similarity, and clusters are subsequently mapped to guideline-derived categories of appropriateness.
Despite its importance, research on NLP-driven analysis of Italian referrals is limited. Venturelli et al. [7] used proprietary commercial software (Clinika VAP [8]) to classify referrals as appropriate or inappropriate, but without explicitly identifying the reasons for prescriptions. To the best of our knowledge, the only other study applying NLP on Italian referrals’ texts is [9], which focused on extracting and analyzing follow-up waiting times from clinical questions rather than assessing referrals appropriateness.
Most previous studies on appropriateness in the Italian NHS relied on structured data or manual text analysis, focusing on medication prescriptions rather than medical examinations [10, 11, 12]. Research on the appropriateness of examinations is more limited [13, 14] and has not incorporated textual data. Other studies focused on the development of decision-support systems for guideline adherence, both in primary care [15, 16] and hospital settings [17, 18]. While effective in specific scenarios [19, 20], such tools are not suited for retrospective analyses of their performance or for quantifying the prevalence of appropriate or inappropriate referrals.
Internationally, several studies have employed NLP and rule-based systems to assess referral appropriateness, often in disease-specific contexts. Haddad et al. [21] analyzed Russian referrals’ texts for appropriateness using labelled data; Sagheb et al. [22] designed a rule-based algorithm to verify guidelines adherence for asthma patients from EHR documents; similarly in [23] a rule-based approach was applied to assess guideline adherence for gout patients. Dimitrov et al. [24] conducted a nationwide evaluation of breast cancer guideline adherence in EHRs using proprietary commercial software (Danny platform [25]), which incorporated rules and manual reviews. Villeana et al. [26] analyzed Chilean referrals’ texts using labelled data to binary classify them for correct priority classes.
However, all these approaches rely on labeled data or proprietary software, limiting their adaptability and scalability across different examination types.
To address these gaps, this study proposes the first fully unsupervised NLP pipeline for analyzing appropriateness in Italian referrals. The pipeline is applicable to any examination for which guidelines define appropriate and inappropriate referral reasons. We developed this approach on a first case study related to venous echocolordoppler of the lower limbs (ECD) and we present results also on a second case study related to flexible endoscope colonoscopy (FEC). Both datasets include all referrals issued during three years in the entire the Lombardy Region, the most populated Italian region, with approximately 10 million inhabitants - akin to a small state.
These two examinations were selected because both are characterized by a high prevalence of inappropriate referrals [27, 28, 29, 30] and had already been targeted by Lombardy’s 2015 resolutions on angiology and gastroenterology [31, 32]. However, no comprehensive study has assessed the impact of those resolutions at scale. Our analysis provides, for the first time, a regional-scale evaluation of referral appropriateness and of guideline effectiveness, which informed the Lombardy Region’s December 2023 resolution [33].
This work thus provides both a methodological contribution, by introducing a scalable NLP pipeline for appropriateness assessment, and an applied contribution, by generating evidence that informed regional policies.
The remainder of this paper is structured as follows: Section 2 describes the datasets and the pipeline, Section 3 presents the results, Section 4 addresses implications, limitations, and future developments; and Section 5 concludes.
2 Methods
2.1 Data
2.1.1 Referrals datasets
The datasets analyzed in this study are extracted from all referrals issued in the Lombardy Region between 2019 and 2021 for ECD and FEC, totalling 496,971 referrals for ECD and 407,949 for FEC. Each referral includes the clinical question, the year, the Local Health Authority (LHA) of the prescribing physician and the type of prescribing physician. Physician types are categorized as general practitioners (GPs) or specialized physicians, while LHAs are the eight healthcare administrative areas within the Lombardy Region. Figure 1 illustrates the distribution of referrals among these groups.
Figure 1: Referrals distributions across years (A, B), type of prescribing physicians (C, D) and LHA (E, F) for ECD (A,C,E) and FEC (B,D,F).
The analysis focuses on the clinical question (CQ), a free-text field typically containing one or a few sentences. These texts exhibit frequent spelling and typing errors, heavy use of abbreviations, inconsistent punctuation, and non-standard capitalization. The dataset for ECD includes 229,509 distinct CQ, with only 15% appearing more than once, while the dataset for FEC includes 89,734 distinct CQ, with 41% appearing more than once. The median length for ECD is 71 characters (IQR=39), while for FEC it is 57 (IQR=20). Only 189 for ECD (0.08%) and 137 for FEC (0.15%) contain a meaningful ICD code, underscoring the necessity of employing NLP for this analysis.
The clinical questions do not contain personal information about patients or physicians.
2.1.2 Manual annotations
To validate the pipeline, a randomly selected subset of 1,000 distinct clinical questions was manually labelled for both ECD and FEC. The labels correspond to categories representing diseases or symptoms that prompted the referral. During an initial review of this subset, ambiguous cases were identified, and an angiologist and a gastroenterologist were consulted to resolve these doubts and align the clusters with prescription guidelines. Annotation guidelines were subsequently refined, and the entire subset was annotated accordingly. Two other reviews of the annotations corrected remaining errors and clarified uncertainties, ensuring the final labelled datasets were error-free. More details on the manual annotations are provided in Additional File 1, Section A.
2.1.3 Prescription guidelines
The guidelines used as a reference to assess referral appropriateness are extracted from [31] for ECD and from [32] for FEC. In particular, we extracted the appropriate and inappropriate diseases and symptoms, considering separately also cases that are more likely to be appropriate or more likely to be inappropriate, depending on details that might not be available in clinical questions. English schematized summaries of these guidelines are reported in Figure 2.
Figure 2: Guidelines for referrals for ECD and FEC in the Lombardy Region. Derived from the the Appendix 3 of [31] and the Appendix 2 of [32], translated into English.
2.2 Evaluation of prescription appropriateness
In this section, we present our pipeline to evaluate prescription appropriateness. The process can be divided into two main steps:
  1. 1.
    Identify the reason for the referral: extract the central semantic meaning from the clinical question.
  2. 2.
    Verify appropriateness: determine if the identified reason aligns with established guidelines.
The absence of labels for this type of data prevented us from adopting a supervised approach. A rule-based approach was excluded since we aim to build a pipeline that can generalize to any examination, and it would not be feasible to develop tailored rules for each type of examination. Moreover, we are interested in understanding the reasons for these referrals, which might include additional and/or more detailed categories than those mentioned in the guidelines. Because of this, the problem was addressed as a clustering problem.
Figure 3 depicts our pipeline with the blocks that compose it, and below, we detail each block.
Figure 3: Schema of the pipeline for assessing referrals reasons and appropriateness
2.2.1 Pre-processing
Physicians write referrals during patients’ examinations, with limited time available. This leads them to write clinical questions that are far from well-readable and correct texts. The grammar of the sentences is often poor, there are frequent typing errors, and the punctuation is either not used or used incorrectly. The use of uppercase letters does not have any particular meaning, and the use of abbreviations is very frequent. Even in the two types of examination analyzed in this study, multiple abbreviations can indicate the same disease/symptoms. In the manually labelled subsets, 15% of clinical questions were marked as having typing errors and 68% as containing abbreviations.
To address these issues, suitable pre-processing steps are applied:
  1. 1.
    lowercasing
  2. 2.
    punctuation removal
  3. 3.
    abbreviation expansion
  4. 4.
    (partial) typos correction
We remove all punctuation and lowercase sentences due to the inconsistent use of these elements.
For abbreviation expansion, we extract all the unique words with length $\leq~3$, and we identify those corresponding to abbreviations of medical terms, defining a vocabulary for their replacement in the texts.
To fix the typing mistakes, all words appearing with frequency $1\%$ and having another word similar to them with frequency $1\%$ are selected. These words are considered misspelt, and the most similar word with frequency $1\%$ is chosen as their correction. The similarity is measured with edit distance, and the minimum threshold is chosen by minimizing errors in a manually reviewed sample of 100 corrections.
Examples of pre-processing effects and an ablation analysis are reported in Additional File 1, Section B.
2.2.2 Text embedding
To appropriately cluster the clinical questions, it is necessary to produce a representation capturing the central semantics of these texts without getting confused by their widely present noise.
Considering the many synonyms and composite terms of the medical jargon used in these texts, vocabulary-based representations, such as TF-IDF [34], cannot be expected to work well for this task.
Static word embeddings, such as Word2Vec [35], can capture the semantic similarities between words with similar meanings (e.g., vasculopathy - arteriopathy, phlebitis - thrombophlebitis) differently from vocabulary-based representations. Nevertheless, they might have more difficulty distinguishing between relevant and irrelevant tokens in our sentences since they do not consider the context in which tokens appear.
On the contrary, context-based embeddings produced by transformers model consider the surrounding context while embedding each token and have often shown superior performances in the recent clinical NLP literature [36, 37, 38, 39]. The problem is particularly challenging in our case due to the limited number of embedding models available for the Italian language, especially for the clinical domain.
To this purpose, we adopt the same model developed in [40] to cluster sentences extracted from Italian discharge summaries, and we verify that even with referrals’ clinical questions, its performances are superior to other models. This model consists of a fine-tuned version of the last Italian version of BERT, Umberto [41], on a dataset of publicly available medical documents in Italian and other languages (translated into Italian) [42].
2.2.3 Clustering
Some studies have shown that HDBSCAN often outperforms other clustering algorithms—such as K-Means, hierarchical clustering, and DBSCAN-when applied to text embeddings, particularly when preceded by a dimensionality reduction step [43, 44]. Text embeddings typically lie on complex, non-linear manifolds and may form clusters with varying densities and irregular shapes. Algorithms like K-Means assume convex cluster geometries and require predefining the number of clusters, which is impractical for our task. Hierarchical clustering is sensitive to noise and outliers and lacks the ability to leave out unclustered data points, often producing forced or less meaningful groupings. In contrast, HDBSCAN does not require the number of clusters to be specified in advance and allows some data points to remain unclustered, making it a better fit for our scenario. Based on these advantages and on a comparison of the results on the ECD dataset, we selected HDBSCAN for the clustering step.
HDBSCAN requires specifying some hyperparameters, including the minimum cluster size, the minimum number of samples in the neighborhood of a core point (min_samplesmin\_samples), and the epsilon. To better accommodate the varying sizes of clusters in our dataset, we adopted a recursive strategy: the minimum cluster size was iteratively reduced, halving at each step, and applied only to data points that remained unclustered in the previous iteration. For the dimensionality reduction, we compared PCA, UMAP and no dimensionality reduction.
We performed a grid search and we selected the configuration with PCA with 80% of variance threshold, reducing the minimum cluster size from 500 to 10, with min_samples=5min\_samples=5 and $\epsilon=0.1$. The complete grid search results, together with comparisons to k-means and agglomerative hierarchical clustering, are reported in Additional File 1, Section C.
Due to resource constraints, it was not feasible to run HDBSCAN on the entire datasets. Therefore, we divided the data into chunks of 30,000 clinical questions and clustered them independently. In later stages, the summarization and mapping procedures allow to merge clusters from different chunks that correspond to the same diseases or symptoms.
2.2.4 Cluster summarization
Summarizing the identified clusters is necessary to understand the symptoms or diseases corresponding to each cluster. To automatize this step, we extracted keywords from each cluster. The most frequent words in each cluster were initially selected, considering a decreasing frequency threshold from 80% until at least two keywords were found. Among these, we excluded those with high frequency in at least half of the clusters. This allowed us to exclude not only stop words but also frequent words for the specific examination we are analyzing.
2.2.5 Mapping clusters to guidelines
After summarization, the keyword representation of the clusters is used to map them to the guidelines. The process involves:
  1. 1.
    Computing embeddings for each clinical question described in the guidelines.
  2. 2.
    Generating an embedding for each cluster based on its keyword summary, enabling direct comparison with the guideline embeddings.
  3. 3.
    Comparing cluster embeddings with guideline embeddings using cosine similarity. Clusters were assigned to the guideline entry with the highest similarity, provided there is one with similarity >0.50>0.50.
2.3 Experiments and Evaluation
In this section, we detail the experimental settings and the evaluation procedure for our pipeline, which are summarized in Figure 4.
Figure 4: Summary of the experiments and evaluation procedure
2.3.1 Experimental Setup and Constraints
All experiments were conducted in an offline environment with limited computational resources: 16 GB of RAM, 16 cores Intel Xeon CPU, no GPU. For security and privacy reasons the data could not be move out of this environment. These constraints, commonly present in healthcare settings, influenced several pipeline design decisions, such as splitting the dataset into chunks for clustering, avoiding re-training Tranformer-based models and avoiding the use of generative large language models.
2.3.2 Evaluation Procedure
The pipeline was evaluated by comparing its output with the labels on the manually annotated subsets. Rather than applying the pipeline solely to the labelled records—which would hinder clustering performance due to limited data—we ran the full pipeline on the entire datasets and then extracted results for the annotated subsets. This ensured meaningful cluster formation and fair evaluation.
We performed evaluation at two levels:
  • Disease label evaluation: We assessed the accuracy of the disease labels produced by the pipeline by comparing them to ground-truth labels. To reduce noise, we retained only labels with more than 10 occurrences (1%) in the labelled subset, grouping all others into an Other category. Clusters not mapped to any label or unclustered clinical questions were assigned to the Other category. Metrics were computed at:
    • Aggregate level: to measure overall precision, recall, and F1-score across all labels.
    • Per-label level: to evaluate performance on individual categories.
  • Appropriateness label evaluation: The quality of the pipeline’s assignment of appropriateness labels was evaluated using a four-class scheme: appropriate, inappropriate, likely (in)appropriate, and undetermined. The likely (in)appropriate category included referrals where reasons mentioned in the guidelines were present, but the text lacked complete information to assess the (in)appropriateness. The undetermined category was used for referrals not addressed in the clinical guidelines. Also in this case, we computed metrics and the aggregate and per-label levels.
At both levels, we computed the metrics at the single clinical questions level and weighting clinical questions by their corresponding referral counts. The latter allows to incorporating the datasets’ real-world distributions into the analysis.
We also qualitatively show t-SNE 2D projections of the embeddings of the labelled clinical questions. Although this type of visualization has known limitations [45], it remains a common approach in the literature for obtaining a qualitative overview of embedding structures [46].
2.3.3 Hyperparameter Selection and Validation Strategy
Clustering algorithm selection and hyperparameter tuning was carried out exclusively on the ECD dataset. The FEC dataset was reserved for external validation only, ensuring that all performance results reported on FEC represent generalization to unseen data and are free from overfitting to the development set.
2.3.4 Comparative Evaluation
To benchmark our pipeline, we compared it with alternative approaches using different text representations and/or clustering methods. These included:
  • Embeddings from the original UmBERTo model (baseline)
  • Embeddings from an Italian Word2Vec model [47]
  • TF-IDF vector representations
  • Latent Dirichlet Allocation (LDA)
  • Direct matching between embeddings of clinical questions and embeddings of diseases from the guidelines
Additional details are provided in Additional File 1, Section C.
2.3.5 Regional and Temporal Analysis
Finally, we applied the pipeline to the full datasets and reported results at a regional level. This provides a complete quantification of referrals for each identified category, together with their appropriateness. Moreover, we present an enrichment analysis stratifying results by year (pre-, during, and post-COVID pandemic), Local Health Authority (LHA), and prescribing physician type (general practitioners vs. specialists).
3 Results
Table 1: Summary evaluation on the disease labels
P-CQ R-CQ F1-CQ P-REF R-REF F1-REF
ECD
Umberto-E3C-Pipeline 86.81 77.32 79.48 92.43 83.28 86.49
Umberto-Base 82.11 73.62 77.21 87.45 77.81 81.89
Word2Vec 76.21 67.55 71.24 81.67 72.35 76.73
TF-IDF 50.11 41.20 42.43 52.54 44.81 45.45
LDA 27.74 34.04 28.48 19.19 27.57 29.90
Embedding Match 68.29 58.44 59.21 70.59 61.52 62.09
FEC
Umberto-E3C-Pipeline 90.42 88.78 88.60 93.59 92.70 92.67
Umberto-Base 84.56 75.24 79.63 88.58 79.66 83.88
Word2Vec 81.21 70.55 75.51 82.45 73.86 77.92
TF-IDF 58.42 51.23 54.59 60.41 53.49 56.74
LDA 35.26 41.25 38.02 31.78 38.64 34.88
Embedding Match 75.69 66.62 70.54 77.11 69.85 73.03
00footnotetext: Summary results of mapping to the disease labels defined by manual annotations on the test sets, computed for clinical questions (CQ) and referrals (REF). P=Precision- R=Recall- F1=F1-Score. The values of all metrics are expressed as percentages.
Table 2: Summary evaluation on the appropriateness labels
P-CQ R-CQ F1-CQ P-REF R-REF F1-REF
ECD
Umberto-E3C-Pipeline 89.24 88.27 88.64 93.58 91.52 92.51
Umberto-Base 86.02 80.78 83.13 89.15 84.97 86.55
Word2Vec 80.66 74.29 77.01 85.41 77.54 80.88
TF-IDF 73.51 68.45 67.92 79.44 75.70 74.22
LDA 47.61 48.43 47.83 42.64 46.31 41.31
Embedding Match 74.82 61.90 62.83 77.52 62.29 64.46
FEC
Umberto-E3C-Pipeline 92.05 90.96 90.88 94.66 93.96 94.05
Umberto-Base 87.41 80.55 83.84 89.62 82.16 85.73
Word2Vec 83.24 74.15 78.43 85.77 75.63 80.38
TF-IDF 67.11 57.96 62.20 68.14 60.85 64.29
LDA 38.96 45.62 42.03 35.41 42.93 38.81
Embedding Match 80.07 69.92 74.20 81.15 73.92 77.15
00footnotetext: Summary results of mapping to the appropriate/inappropriate/likely (in)appropriate/undetermined categories on the manually labelled test sets, computed for clinical questions (CQ) and referrals (REF). P=Precision- R=Recall- F1=F1-Score. The values of all metrics are expressed as percentages.
3.1 Pipeline evaluation
Table 1 presents the summary results of the evaluation performed on the manually annotated subsets, at the disease level, comparing different approaches. The pipeline with our fine-tuned Umberto embeddings demonstrates superior performance compared to alternative text representations and clustering algorithms, confirming the importance of contextual representations. The approach with direct matching between embeddings of clinical questions and embeddings of diseases from guidelines exhibits discrete results, but lower than our pipeline, confirming that the clustering procedure act as a useful smoothing step. The metrics are computed for both clinical questions (CQ) and referrals (REF). The referral-level scores are higher for all methods, except LDA, due to the influence of clinical questions that occur more frequently, which seem to be easier to cluster for them, being less detailed.
Table 3: Detailed results of the cluster mapping to the appropriateness labels for our pipeline
P-CQ R-CQ F1-CQ #CQ P-REF R-REF F1-REF # REF
ECD
Appropriate 98.55 93.15 95.77 365 98.97 93.95 96.39 512
Inappropriate 91.29 81.88 86.33 320 96.49 89.58 92.91 921
Likely inappr. 43.28 89.23 58.29 65 84.03 89.89 90.86 628
Undetermined 84.62 79.20 81.82 250 87.22 83.33 85.23 426
FEC
Appropriate 99.36 89.44 94.14 521 99.88 96.96 98.40 2501
Inappropriate 95.35 87.23 91.11 47 95.71 88.16 91.78 76
Likely appr. 98.27 87.18 92.39 195 98.88 90.72 94.62 388
Undetermined 75.24 100.00 85.87 237 84.19 100.00 91.42 591
00footnotetext: Detailed results of the cluster mapping to the appropriateness labels for our pipeline, on the manually labelled test sets, computed for clinical questions (CQ) and referrals (REF). P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
Table 2 shows the same results achieved by grouping labels into the appropriate, inappropriate, likely (in)appropriate and undetermined categories. These results show higher aggregate performance scores for ECD, compared to Table 1, indicating that most errors occurred within the same appropriateness category rather than across categories. The results for FEC are also superior, but the improvement is limited. This suggests that distinguishing between appropriate and inappropriate referrals is simpler, while finer distinctions at the label level are more challenging. Table 3 details the metric by appropriateness label, for our pipeline. At the clinical question level, it highlights issues with the likely inappropriate class for ECD, which corresponds directly to the varices class at disease level, and it shows slightly lower results for the undetermined class, but at the referrals level the only relevant difference is with the undetermined class for ECD.
Table 4: Detailed results of the cluster mapping for each disease label for our pipeline on ECD dataset
Category P-CQ R-CQ F1-CQ # CQ P-REF R-REF F1-REF # REF
Claudicatio 83.33 45.45 58.82 11 90.00 56.25 69.23 16
Pain 85.00 91.07 87.93 56 88.61 93.33 90.91 75
Edema 85.47 96.15 90.50 104 89.51 97.32 93.25 149
Venous insuff. 94.90 74.50 83.47 200 98.06 88.09 92.81 747
Paresthesis 100.00 81.25 89.66 16 100.00 88.00 93.62 25
Saphenectomy 100.00 92.59 96.15 27 100.00 93.94 96.88 33
(Trombo) Phlebitis 98.70 93.83 96.20 81 99.09 93.16 96.04 117
Thrombosis 96.24 94.04 95.13 218 96.99 94.77 95.87 306
Ulcer 95.83 92.00 93.88 25 97.44 92.68 95.00 41
Varices 43.28 89.23 58.29 65 84.03 98.89 90.86 628
Varices surgery 100.00 42.86 60.00 14 100.00 46.67 63.64 15
Vasculopathy 82.14 44.23 57.50 52 95.33 68.92 80.00 148
Other 63.57 67.94 65.68 131 62.56 70.59 66.33 187
00footnotetext: Detailed results of the cluster mapping for each disease for our pipeline, on the manually labelled test set for ECD, computed for clinical questions (CQ) and referrals (REF). Labels with support <10<10 are grouped with Other. P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
Table 5: Detailed results of the cluster mapping for each disease label for our pipeline on FEC dataset
Category P-CQ R-CQ F1-CQ # CQ P-REF R-RE F1-REF # REF
Bleeding 100.00 77.91 87.59 163 100.00 97.31 98.64 1748
Polyps 99.35 97.47 98.40 158 99.71 98.55 99.12 344
Family history / prevention 97.56 86.33 91.60 139 98.57 91.39 94.85 302
Colon cancer 96.84 85.98 91.09 107 98.24 89.30 93.56 187
Abdominal pain 88.57 89.42 89.00 104 93.66 89.72 91.65 214
Anaemia 92.75 98.46 95.52 65 95.73 97.39 96.55 115
Chronic inflammatory bowel disease 100.00 89.29 94.34 56 100.00 88.37 93.83 86
Constipation 78.38 80.56 79.45 36 91.61 94.24 92.91 139
Diverticulosis 90.00 81.82 85.71 33 86.67 81.25 83.87 48
Altered bowel 84.85 93.33 88.89 30 92.54 95.38 93.94 65
Chronic diarrhoea 100.00 82.14 90.20 28 100.00 93.46 96.62 107
Diverticulitis 100.00 92.86 96.30 14 100.00 89.29 94.34 28
Other 47.14 98.51 63.77 67 59.93 99.42 74.78 173
00footnotetext: Detailed results of the cluster mapping for each disease for our pipeline, on the manually labelled test set for FEC, computed for clinical questions (CQ) and referrals (REF). Labels with support <10<10 are grouped with Other. P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
Tables 4 and 5 detail metrics by disease for our pipeline. Most diseases achieve F1-CQ $80\%$, with many exceeding $90\%$. Lower scores are observed for some ECD categories such as varices, varices surgery, vasculopathy, claudicatio and other, while for FEC only other exhibits poor precision. At the referral level (F1-REF), only claudicatio, varices surgery and other achieve scores below $80\%$ for ECD, while for FEC only other is below $80\%$.
Figure 5: Confusion matrix of the clustering assignment to the labels on the manually annotated test sets, for ECD (A, B) and FEC (C,D) computed considering the number of clinical questions (A, C) and the number of referrals (B, D)
The confusion matrices shown in Figure 5 allow for a more detailed analysis of misclassification errors. For ECD, errors in varices surgery are often due to misclassifications as varices, a strictly related cluster. Regarding varices, its precision is mainly affected by misclassifications from venous insufficiency and other. Looking at the specific clinical questions manually labelled as venous insufficiency and classified as varices, we notice that some present both diagnoses, making this misclassification reasonable. For the remaining ECD classes, misclassified samples fall mainly into the other group, affecting its precision. For FEC, the majority of misclassification errors are wrong assignments to the other class. A few are wrong assignments to abdominal pain, which are in multiple cases due to the mention of this symptom along with others that were more relevant. A detailed list of all identified clusters, together with their mapping, is reported in Additional File 1, Section C.
Figure 6 displays t-SNE 2D projections of the embedding vectors of the clinical questions in the two labelled subsets, showing that for most classes there are one or more distinct regions of the 2D space that tend to concentrate the corresponding clinical questions.
Figure 6: 2D t-SNE projections of embeddings of clinical questions from the validation sets of ECD (A) and FEC (B) with their labels, excluding the Other class
3.2 Analysis of the entire dataset
Table 6: Summary of referrals’ clusters for ECD with appropriateness labels
Cluster Label # REF %
[…]
00footnotetext: Summary of referrals’ clusters for ECD, mapped to all labels defined in manual annotations. Green indicates appropriate clusters, red inappropriate clusters, yellow clusters for which the information might not be sufficient to define the appropriateness but is more likely inappropriate, and light grey those not mentioned in the guidelines.
Table 7: Summary of referrals’ clusters for FEC with appropriateness labels
Cluster Label # REF %
[…]
00footnotetext: Summary of referrals’ clusters for FEC, mapped to all labels defined in manual annotations. Green indicates appropriate clusters, red inappropriate clusters, yellow clusters for which the information might not be sufficient to define the appropriateness but is more likely appropriate, and light grey those not mentioned in the guidelines.
Tables 3.2 and 1 extend the analysis to the entire ECD and FEC datasets, highlighting the appropriateness labels for each cluster. We can observe that the distribution of the reasons for the referrals is unbalanced, with 75% of referrals covered by the first five reasons in ECD and by the first seven reasons in FEC. In ECD, varices cluster was assigned to likely inappropriate since this diagnosis is inappropriate unless it is associated with surgery, an element that is often difficult to exclude, even for a human reader. In FEC, colon cancer familiarity / prevention and chronic inflammatory bowel disease were instead marked as likely appropriate since guidelines state that physicians should carefully consider the patient history and condition to establish if and when they should be performed. Specific conditions mentioned in the guidelines — such as requiring ten years from diagnosis or significant colon involvement — were largely absent from referral texts.
Summarizing Table 3.2, for ECD, 34.32% of referrals appear to be appropriate, while 34.07% seems to be not appropriate. The remaining 31.61% referrals can be further divided into 14.37% more likely to be not appropriate, 10.20% that are not mentioned on the guidelines, and 7.04% that cannot be clustered or cannot be mapped to the labels defined by manual annotations. For FEC, 55.98% of referrals appear to be appropriate, while only 5.33% seems to be not appropriate. There are then 15.95% more likely to be appropriate, 15.77% that are not mentioned on the guidelines, and 6.98% that cannot be clustered or cannot be mapped to the labels defined by manual annotations. Notably, in both datasets the other group is limited, while there are clusters for many symptoms and diseases not mentioned in the guidelines, especially for ECD. Although individually rare, these uncovered reasons accounted for a non-negligible proportion of overall referrals, suggesting the need for a broader guideline scope. Additionally, we observed ambiguity in referrals citing varices, where physicians often omitted whether surgical intervention was required. In contrast, FEC referrals showed higher levels of appropriateness, though some inappropriate clusters were identified.
Figure 7 provides some examples, along with their English translation, of clustered clinical questions.
Figure 7: Examples of clustered clinical questions. English translations do not completely reflect the abbreviations and typos of the original Italian texts.
Figure 8: Prescription appropriateness stratified by LHA (A, B), type of prescribing physicians (C, D) and year (E, F) for ECD (A,C,E) and FEC (B,D,F).
Figure 8 reports the level of appropriateness stratified by LHA, type of prescribing physicians and year, providing additional insights into the reasons for inappropriate prescriptions. This stratified analysis reveals significant geographical variability for both examinations and especially for ECD, suggesting that targeted interventions are needed in specific Local Health Authorities. General practitioners exhibited higher rates of inappropriate ECD referrals compared to specialist physicians. This might have multiple reasons, including a larger patient base, more frequent encounters with less severe cases, and a lower knowledge and/or agreement with guidelines [48, 49]. For FEC, the rate of inappropriateness is similar but specialized physicians have more appropriate referrals. An important trend emerges when analyzing the impact of COVID-19 on referral practices. During the early pandemic phase in 2020, inappropriate ECD referrals dropped significantly, likely due to resource constraints prompting physicians to prioritize urgent cases. This pattern is consistent with existing literature, which reports a general reduction in non-urgent medical procedures and a shift toward more conservative referral behaviour during the pandemic [50, 51, 52, 53]. However, this decline was temporary, with inappropriate referrals rebounding in 2021 as healthcare services normalized. A similar, though less pronounced, trend was observed for FEC, possibly due to the already limited number of inappropriate referrals for this examination.
4 Discussion
In this study, we proposed a methodological pipeline leveraging NLP techniques to assess the appropriateness of diagnostic referrals in the Italian NHS. We illstrated its use through two case studies on venous echocolordoppler of the lower limbs in angiology and flexible endoscope colonoscopy in gastroenterology. By applying an unsupervised approach, we uncovered key trends in referral motivations and guideline adherence, offering the first large-scale assessment of this kind in Italy.
Our work presents several strengths. The pipeline operates in an unsupervised setting, without the need for labelled data, which are costly and time-consuming to produce. This makes it applicable to any examination with established guidelines defined in terms of appropriate and inappropriate reasons for referrals. It is also designed to run on large datasets without requiring high computational resources, lowering the barrier for adoption. In addition to quantifying appropriateness, the pipeline offers a detailed picture of referral motivations, thereby revealing reasons not explicitly covered by existing guidelines.
We demonstrated these capabilities on two population-scale datasets, covering an entire Italian region with 10 million inhabitants for three years. Validation on manually annotated subsets confirmed the robustness of our pipeline and demonstrated that contextual embedding models significantly outperform simpler clustering techniques in capturing the semantic nuances of referral texts. We observed consistently high performances at the appropriate/inappropriate level and for most diseases/symptoms, with only two clusters in ECD with lower recall and one cluster in FEC with lower precision. The excellent results on two different examinations, one of which was not used during pipeline development, confirmed the generalizability of our methodology.
The large-scale analysis enabled, for the first time, a comprehensive view of referral motivations for these examinations. This is a relevant result for public health authorities, given the absence of structured recording of this information in the Italian NHS. The identification of undocumented reasons for referrals highlighted potential gaps in the existing guidelines and the stratified analyses further revealed geographical and temporal patterns and variations across physician types, consistent with findings in the literature and suggestive of targeted interventions to reduce inappropriate referrals. Importantly, the findings directly supported the Lombardy Region in issuing a new resolution extending and reinforcing referral guidelines. This represents one of the first documented cases in Italy where an AI-based analysis of routine healthcare data directly shaped health policy.
This study is a relevant example of a real-world, large-scale application of artificial intelligence in healthcare. Despite a growing body of research, the adoption of AI-based solutions in clinical practice faces significant challenges [54] and remains largely dominated by diagnostic tools in medical imaging [55, 56], particularly in Italy [57]. Our work points to a complementary direction: leveraging large-scale textual data to monitor referral practices. In this way, it illustrates how AI can extend its role beyond diagnosis to support healthcare governance and inform evidence-based policy.
Despite its strengths, this study has some limitations. Manual annotation was limited to a subset of the data due to time constraints, and performed by a single annotator, limiting the possibility of inter-annotator agreement analysis. Our approach also assigns a single primary reason per referral, potentially overlooking multi-reason referrals, which could be better addressed through fuzzy clustering techniques. Moreover, the lack of access to patient histories limited our ability to fully assess clinical appropriateness in certain cases, such as referrals for varices, where prior surgical indication is relevant. It is important to clarify that our pipeline does not aim to assess the appropriateness of individual referrals. Instead, our approach provides a systematic, automated, and large-scale evaluation of prescription practices at the regional level, offering valuable insights for policymakers and healthcare administrators.
In future work, we aim to extend the validation of the pipeline to other examinations and broader settings, including larger annotated datasets and other Italian regions, to further support the adoption of data-driven tools for evaluating referral practices and improving healthcare delivery.
5 Conclusions
This study proposed a methodological pipeline for assessing the level of appropriateness of Italian referrals. We presented two case studies on venous echocolordoppler of the lower limbs and flexible endoscope colonoscopy in the Lombardy Region, demonstrating excellent performance on manually annotated test sets and deriving relevant insights for public health authorities. The outcome of this analysis allowed the Lombardy Region to reinforce existing guidelines for physicians, with the aim of reducing the number of inappropriate referrals. This work is an example of how artificial intelligence can be deployed on large, real-world administrative datasets, highlighting the potential of AI to support healthcare systems in monitoring practices, informing policy, and ultimately alleviating significant burdens on the Italian NHS by lowering waiting times and enhancing overall effectiveness of care.
Format: pdf Title: Supplementary Material Description: contains more details about the manual annotation process, hyperparameters selection and additional results
[…]
List of abbreviations
  1. 1.
    AI: Artificial Intelligence
  2. 2.
    BERT: Bidirectional Encoder Representation for Transformers
  3. 3.
    CIBD: Chronic Inflammatory Bowel Disease
  4. 4.
    CQ: Clinical Question
  5. 5.
    DGR: Decreto della Giunta Regionale (Regional Law)
  6. 6.
    ECD: venous EcoColorDoppler of the lower limbs
  7. 7.
    EHR: Electronic Health Record
  8. 8.
    FEC: Flexible Endoscope Colonscopy with possible biopsies
  9. 9.
    GP: General Practitioner
  10. 10.
    ICD: International Classification of Diseases
  11. 11.
    IQR: InterQuartile Range
  12. 12.
    LDA: Latent Dirichlet Allocation
  13. 13.
    ATS/LHA: Agenzia di Tutela della Salute / Local Health Authority
  14. 14.
    NHS: National Health Service
  15. 15.
    NLP: Natural Language Processing
  16. 16.
    REF: Referral
  17. 17.
    TF-IDF: Term Frequency - Inverse Document Frequency
Declarations
Ethics approval and Consent to participate
The need for ethics approval is deemed unnecessary according to national regulations. In particular, according to the rules from the Italian Medicines Agency (available at: https://www.gazzettaufficiale.it/eli/id/2024/08/20/24A04320/sg), retrospective studies using administrative databases do not require Ethics Committee protocol approval.
The need for consent to participate is deemed unnecessary according to national regulations. In particular, see the rules of the Italian Privacy Authority available at https://www.garanteprivacy.it/web/guest/home/docweb/-/docweb-display/docweb/9124510 and the European Union regulation n. 2016/679 (GDPR) Art. 1 c. 1 (https://gdpr.eu/article-1-subject-matter-and-objectives-overview/), Art. 4 c.1 (https://gdpr.eu/article-4-definitions/) and Recital 26 (https://gdpr-info.eu/recitals/no-26/).
Consent for publication
Not applicable.
Availability of data and materials
The data that support the findings of this study are available from Regione Lombardia but restrictions apply to the availability of these data, which were used under license for the current study, and so are not publicly available. Data are however available from the authors upon reasonable request and with permission of Regione Lombardia.
Code availability
Competing interests
The authors declare that they have no competing interests.
Funding
No fundings supported this research.
Authors’ contributions
VT: analysis and interpretation of data, development of the methodology, drafting of the manuscript AB: concept and design of the study, acquisition, analysis and interpretation of data ME: concept and design of the study, acquisition, analysis and interpretation of data OL: concept and design of the study, acquisition, analysis and interpretation of data FI: concept and design of the study, analysis and interpretation of data, critical revision of the manuscript, supervision. All authors read and approved the final manuscript.
Acknowledgements
The present research is part of the activities of “Dipartimento di Eccellenza 2023-2027”. We thank the Welfare Directorate of Regione Lombardia (Agreement between Politecnico di Milano and Regione Lombardia DGR n. XI/5301-27/09/2021) and, in particular, the working group on waiting list monitoring. We thank Dr. Paola Colombo for the consultation regarding the reasons for angiology referrals and Dr. Marco Soncini for gastroenterology referrals.
Appendix A Manual annotations
A validation set of 1,000 clinical questions was manually annotated for the main disease/symptom determining the referrals, for both ECD and FEC datasets. The existing guidelines from the Lombardy Region [31, 32] were used as a reference, but labels were not forced to be one of the categories mentioned in the guidelines as appropriate or inappropriate. If none of the categories mentioned in the guidelines fit the reason for the referral, a new category representing the main disease or symptom mentioned in the referral was created. The other category was used only when it was impossible to identify a symptom or a disease determining the referral (e.g., “in-depth examination to actually establish whether there is an ongoing pathology given the symptoms present” or “gfwhdfg”).
In case of multiple diseases or symptoms, the following priority order was defined:
  1. 1.
    Reasons mentioned as appropriate in the guidelines
  2. 2.
    Reasons mentioned as inappropriate among the guidelines
  3. 3.
    Reasons not mentioned in the guidelines
If considering these priority levels, there were still multiple diseases/symptoms, the following priority order was defined:
  1. 1.
    If there is one symptom/disease which looks like the leading cause, prefer it
  2. 2.
    If there are multiple symptoms/diseases mentioned in the guidelines, choose the first one to be mentioned in the text
  3. 3.
    If there are multiple symptoms/diseases not mentioned in the guidelines, prefer those for which a category was already created, otherwise the first one to be mentioned in the text
This procedure, after four iterations, including consultation with specialist physicians, defined the labels reported in Tables A.1 and A.2. After review of these labels, some of them were merged:
  • bleeding and occult blood since bleeding in this context is always an occult bleeding
  • colitis, chron’s disease, ulcerative rectocolitis, chronic inflammatory bowel disease, under the latter which is a category comprehensive of all these diseases
Table A.1: Labels derived by manual annotations on the validation set of 1,000 clinical questions for ECD.
Label # CQ # REF
Thrombosis 218 306
Venous insufficiency 200 747
Edema 104 149
(Thrombo)phlebitis 81 117
Varices 65 628
Pain 56 75
Vasculopaty 52 148
Safenectomy 27 33
Surgery 26 26
Ulcer 25 41
Paresthesis 16 25
Varices surgery 14 15
Other 13 16
Claudicatio 11 16
Swelling 9 10
Transplanctation 8 8
Teleangectasias 7 7
Cardiopathy 6 15
Dermatitis 6 12
Dimer 6 6
Ischaemia 5 5
Diabetes 4 5
Prosthesis 4 4
Cancer 3 5
Heavy Legs 3 4
Hypertension 3 6
Stent 3 2
Arthritis 2 3
Calcification 2 2
Cold legs 2 2
Dyslipidaemia 2 2
Tingling 2 3
Dialysis 1 1
Dyspnoea 1 1
Embolism 1 2
Haemorrhage 1 1
Hypercholesterolaemia 1 3
Hypertrophy 1 1
Hyposthesia 1 1
Infection 1 1
Lymphangitis 1 1
Metabolic syndrome 1 1
Multiple sclerosis 1 1
Neuropathy 1 1
Pregnancy 1 1
Smoker 1 1
Vascular encephalopathy 1 1
Table A.2: Labels derived by manual annotations on the validation set of 1,000 clinical questions for FEC
Label # CQ # REF
Polyps 158 344
Bleeding 127 959
Colon cancer 107 187
Abdominalgia 104 214
Colon cancer familiarity 82 141
Anaemia 65 115
Prevention 57 161
Occult blood 36 789
Constipation 36 139
Diverticulosis 33 48
Altered bowel 30 65
Colitis 28 40
Chronic diarrhoea 28 107
Diverticulitis 14 28
Crohn’s disease 12 16
Haemorrhoidal disease 9 30
Chronic inflammatory bowel disease 8 15
Ulcerative rectocolitis 8 15
Surgery 5 6
Intestinal obstruction 5 10
Dyspepsia 4 5
Weight loss 3 8
Irritable bowel syndrome 3 6
Lymphoma 3 4
Transplant 3 15
Other 2 2
Colonpathy 2 7
Enteritis 2 3
Proctitis 2 4
Rectal prolapse 2 12
Lynch syndrome 2 2
Angiodysplasia 1 1
Appendicitis 1 1
Celiac disease 1 1
Intestinal loops distension 1 1
Hepatopathy 1 1
Protruding formation 1 1
Gastritis 1 2
Ileitis 1 1
Hypercholesterolaemia 1 1
Splenic flexure thickening 1 3
Rectal wall thickening 1 2
Sigma wall thickening 1 1
Behcet’s disease 1 1
Liver disease 1 1
Pancreatitis 1 1
Proctalgia 1 3
Sigmoiditis 1 1
Suspected colic lesion 1 33
Splenic fissure stenosis 1 1
Rectal tenesmus 1 2
Appendix B Pre-processing results
Table B.1 reports the list of the 50 most frequent abbreviations that were expanded by our pre-processing, while Table B.2 contains an example list of 50 misspelt words that were detected and corrected during our pre-processing. Table B.3 reports the summary results on disease labels without applying pre-processing.
Table B.1: Most frequent abbreviations in clinical questions for echocolordoppler of the venous limbs, along with their expansions in Italian and English.
Abbr Freq Expansion Expansion [ENG]
dx 41942 destra right
sx 35400 sinistra left
tvp 35392 trombosi venosa profonda deep vein thrombosis
inf 26644 inferiore inferior
sin 8490 sinistra left
pz 6100 paziente patient
ndd 4555 di natura da determinare of a nature to be determined
$a$ 4079 arti limbs
tvs 3650 trombosi venosa superficiale superficial venous thrombosis
sn 3055 sinistra left
x 2210 per for
k 2170 cancro cancer
ivc 2124 insufficienza venosa cronica chronic venous insufficiency
$i$ 1768 inferiori inferior
tep 1695 tromboembolia polmonare pulmonary thromboembolism
nao 1291 nuovi anticoagulanti orali new oral anticoagulants
ca 1262 cancro cancer
ds 1001 destra right
paz 959 paziente patient
tao 924 terapia anticoagulante orale oral anticoagulant therapy
vgs 911 vena grande safena great saphenous vein
tv 875 trombosi venosa venous thrombosis
tp 852 trombosi profonda deep thrombosis
s 819 sindrome syndrome
dm 800 diabete mellito diabetes mellitus
bil 792 bilaterale bilateral
irc 650 insufficienza renale cronica chronic renal failure
ins 620 insufficienza failure
sup 608 superiore upper
fa 592 fibrillazione atriale atrial fibrillation
v 576 venosa venous
iv 497 insufficienza venosa venous insufficiency
spt 469 sindrome post trombotica post-thrombotic syndrome
npl 451 neoplasia neoplasm
iii 444 inferiori inferior
ep 421 episodio episode
tx 364 trapianto transplantation
ctr 351 controllo control
art 327 tromboembolismo venoso venous thromboembolism
tpv 326 trombosi profonda venosa deep vein thrombosis
vs 307 vena safena saphenous vein
cr 297 cronica chronic
$g$ 269 giorni days
tev 256 tromboembolia venosa venous thromboembolism
pta 251 angioplastica percutanea transluminale percutaneous transluminal angioplasty
ven 235 venosa venous
int 227 intervento surgery
mvc 225 malattia venosa cronica chronic venous disease
ev 220 eventuale possible
$t$ 203 tibio-tarsica tibio-tarsal
Table B.2: Sample of 50 misspelt words with the corresponding replacements applied by our pre-processing.
Original Word Corrected Word Corrected Word [ENG]
iprtensioone ipertensione hypertension
stwnosi stenosi stenosis
areomasica ateromasia atheromasia
aduta caduta fall
clacifiche calcifiche calcifications
lipitimiche lipotimie lipothymias
esignificative significative significant
veriginosa vertiginosa vertiginous
lipotimoco lipotimici lipotimics
emicrani emicrania migraine
cardiopatiia cardiopatia heart disease
30anni anni years
sinndrome sindrome syndrome
vertigicnosa vertiginosa vertiginous
eriosa arteriosa arterial
srteriopatia arteriopatia arteriopathy
carotgidee carotidea carotid
nvertiginosa vertiginosa vertiginous
nlipotimia lipotimici lipothymic
pregressotia pregresso previous
priomo primo first
dislipemua dislipidemia dyslipidaemia
caroltidea carotidea carotid
iportesa ipertesa hypertensive
cerebricolare cerebrovascolare cerebrovascular
neurologoin neurologo neurologist
cronarico coronarica coronary
arteiopatia arteriopatia arteriopathy
bilaterati bilaterale bilateral
aritimica aritmica arrhythmic
bpcoe bpco bpco
extrasisoliche extrasistolia extrasystole
paiente paziente patient
iprtnsione ipertensione hypertension
teriopati arteriopatia arteriopathy
caprtidea carotidea carotid
soszpetta sospetta suspected
dislidemica dislipidemia dyslipidaemia
lipotimicio lipotimici lipothymics
epeisodi episodi episodes
ipertrigliceridiemia ipertrigliceridemia hypertriglyceridaemia
ertiginosa vertiginosa vertigo
iperpensione ipertensione hypertension
neopla neoplasia neoplasia
oderata moderata moderate
cardiopataia cardiopatia cardiopathy
polidistrettuae polidistrettuale polydistrict
dislipidemiaecocolo dislipidemico dyslipidaemic
ipwrtensone ipertensione hypertension
acugene acufene tinnitus
Table B.3: Results of clusters mapped to the disease labels on the manually labelled test sets, computed for clinical question (CQ) and referrals (REF) without applying pre-processing steps. P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
P-CQ R-CQ F1-CQ P-REF R-REF F1-REF
ECD
Umberto-E3C-Pipeline 84.01 73.51 78.40 89.15 79.28 83.95
Umberto-Base 77.99 70.68 74.18 85.01 73.18 78.66
Word2Vec 75.63 62.25 68.27 78.90 67.77 72.99
TF-IDF 46.28 33.61 38.92 45.73 37.83 41.46
LDA 24.38 32.27 27.88 16.51 25.97 20.24
FEC
Umberto-E3C-Pipeline 88.59 85.53 87.12 90.99 87.65 89.32
Umberto-Base 80.28 70.31 75.02 85.79 75.42 80.35
Word2Vec 78.66 68.55 73.27 80.15 70.38 74.96
TF-IDF 55.45 44.91 49.69 55.95 49.60 52.65
LDA 31.52 36.30 33.80 26.70 33.91 29.86
Appendix C Clustering
C.1 Hyperparameters and algorithms selection
Table LABEL:tab:hdbscan_hyper reports the grid search for HDBSCAN hyperparameters, evaluated with F1-CQ on ECD dataset.
Table C.1: Grid search for HDBSCAN hyperparameters, with F1-CQ on ECD dataset
min_cluster_size min_samples cluster_selection_epsilon F1-CQ [%]
20 0 11.42
20 0.1 11.42
20 0.5 11.42
20 1 0 63.91
20 1 0.1 63.91
20 1 0.5 63.91
20 5 0 10.13
20 5 0.1 10.13
20 5 0.5 10.13
20 10 0 77.85
20 10 0.1 77.85
20 10 0.5 77.85
20 20 0 11.42
20 20 0.1 11.42
20 20 0.5 11.42
50 0 40.31
50 0.1 40.31
50 0.5 40.31
50 1 0 60.81
50 1 0.1 60.81
50 1 0.5 60.81
50 5 0 71.72
50 5 0.1 71.72
50 5 0.5 71.72
50 10 0 70.40
50 10 0.1 70.40
50 10 0.5 70.40
50 20 0 40.31
50 20 0.1 40.31
50 20 0.5 40.31
100 0 23.87
100 0.1 23.87
100 0.5 23.87
100 1 0 75.64
100 1 0.1 75.64
100 1 0.5 75.64
100 5 0 44.60
100 5 0.1 44.60
100 5 0.5 44.60
100 10 0 51.82
100 10 0.1 51.82
100 10 0.5 51.82
100 20 0 40.31
100 20 0.1 40.31
100 20 0.5 40.31
200 0 14.46
200 0.1 14.46
200 0.5 14.46
200 1 0 67.41
200 1 0.1 67.41
200 1 0.5 67.41
200 5 0 36.43
200 5 0.1 36.43
200 5 0.5 36.43
200 10 0 43.32
200 10 0.1 43.32
200 10 0.5 43.32
200 20 0 31.80
200 20 0.1 31.80
200 20 0.5 31.80
500 0 18.64
500 0.1 18.64
500 0.5 18.64
500 1 0 15.48
500 1 0.1 15.48
500 1 0.5 15.48
500 5 0 14.45
500 5 0.1 14.45
500 5 0.5 14.45
500 10 0 23.37
500 10 0.1 23.37
500 10 0.5 23.37
500 20 0 16.31
500 20 0.1 16.31
500 20 0.5 16.31
step from 500 0 34.52
step from 500 0.1 34.52
step from 500 0.5 34.52
step from 500 1 0 71.54
step from 500 1 0.1 71.54
step from 500 1 0.5 71.54
step from 500 5 0 79.48
step from 500 5 0.1 79.48
step from 500 5 0.5 79.48
step from 500 10 0 63.28
step from 500 10 0.1 63.28
step from 500 10 0.5 63.28
step from 500 20 0 41.95
step from 500 20 0.1 41.95
step from 500 20 0.5 41.95
Table C.1: *
Figures C.1 and C.2 report the performance of K-Means and Agglomerative Hierarchical Clustering on the ECD dataset, at the CQ level, varying the number of clusters. Their results are consistently lower than those from HDBSCAN.
Figure C.1: K-means performance on ECD-CQ varying the number of clusters
Figure C.2: Agglomerative hierarchical clustering performance on ECD-CQ varying the number of clusters
For TF-IDF representation, we limited the vocabulary to the top 1,000 most frequent tokens. The best results presented are obtained with K-Means. The number of clusters was selected with the knee-elbow method analyzing WSS with k from 2 to 50, which led to k=40 for ECD and k=35 for FEC.
For LDA, the number of topics (clusters) was selected by analyzing coherence with k from 2 to 50, which led to k=7 for ECD and k=9 for FEC.
For direct embeddings match, we matched Umberto embeddings of clinical questions with embeddings of known diseases/symptoms from guidelines, using cosine similarity. If no element in the guidelines matched with a similarity of at least 0.6 we assigned the Other label.
C.2 Additional clustering results
Tables LABEL:tab:clusters-summ and LABEL:tab:clusters-summ-fec present all the clusters identified for ECD and FEC, respectively, with their keyword descriptions, the sizes and the labels they were assigned.
Table C.2: Summary of all clusters identified for ECD, including their keyword representation, the label assigned them and the number of distinct clinical questions (CQ) and referrals (REF) that they contain.
ID CLUSTER KEYWORDS [ITA] CLUSTER KEYWORDS [ENG] CLUSTER LABEL # CQ # REF
1 claudicatio; riferita claudication; referred claudication 1132 1686
2 claudicatio; intermittens claudicatio; intermittens claudication 653 1179
3 claudicatio claudicatio claudication 402 530
4 claudicatio; sospetta; claudicazio claudicatio; suspected; claudication claudication 60 126
5 claudicatio; arti; inferiori claudicatio; limbs; lower claudication 41 49
6 dermatite dermatitis dermatitis 421 474
7 diabete diabetes diabetes 1540 3249
8 diabete; mellito diabetes; mellitus diabetes 534 1847
9 diabete; cardiopatia; mellito; diabetes; heart disease; mellitus diabetes 755 1254
10 diabetico; diabetica diabetes; diabetic diabetes 209 165
11 dislipidemia; diabete dyslipidaemia; diabetes diabetes 36 68
12 dialisi dialysis dialysis 110 222
13 dimero dimer dimer 1343 1601
14 dispnea dyspnoea dyspnoea 104 171
15 dolore pain pain 6343 12051
16 deambulazione; dolore; difficolta walking; pain; difficulty pain 1294 3884
17 algia; artralgia; piedi algia; arthralgia; foot pain 1028 2491
18 polpaccio; dolore calf; pain pain 707 1753
19 algie; arti; gambe; polpacci; bilaterali; natura; determinare pain; limbs; calves; bilateral; nature; determine pain 160 631
20 dolore; natura; determinare; algie pain; nature; determine; pain pain 53 126
21 dolore; popliteo pain; popliteal pain 35 117
22 algia; covid; gamba pain; covid; leg pain 30 63
23 dolore; polpaccio; natura; determinare pain; calf; nature; determine pain 12 10
24 crampi; notturni cramps; nocturnal pain 6 5
25 edema edema edema 11686 20723
26 edemi; declivi edema; declivities edema 9491 16473
27 edema; arti; edemi edema; limbs; edemas edema 3712 7867
28 arti; inferiori; edemi limbs; lower; edemas edema 528 2486
29 edema; gamba; piede; polpaccio edema; leg; foot; calf edema 1046 2467
30 edema; arti; inferiori edema; limbs; lower edema 899 2350
31 edema; perimalleolare edema; perimalleolar edema 1039 2214
32 edema; natura; determinare edema; nature; determine edema 1027 2199
33 edemi; perimalleolari edema; perimalleolar edema 1190 2151
34 edema; malleolare edema; malleolar edema 935 1136
35 edemi; vasculopatia; declivi; edema edema; vasculopathy; declivities edema 529 602
36 edema; flebite edema; phlebitis edema 217 297
37 arti; bilaterale; piede; edema; succulenza; perimalleolare limbs; bilateral; foot; edema; succulence; perimalleolar edema 117 192
38 arti; inferiori; edema limbs; lower; edema edema 80 168
39 arti; inferiori; natura; determinare; edema; edemi limbs; lower; nature; determine; edema edema 80 168
40 edema; inferiori; arti edema; lower; limbs edema 74 163
41 edema; sinistra edema; left edema 61 108
42 claudicatio; edemi; edema; cardiopatia; arti; inferiore; lieve claudication; edema; edema; cardiopathy; limbs; lower; mild edema 32 34
43 linfedema; arti lymphedema; limbs edema 23 24
44 emorragia haemorrhage haemorrhage 18 29
45 formicolio tingling tingling 43 170
46 freddo cold cold 49 147
47 fumatore smoker smoker 121 193
48 gravidanza pregnancy pregnancy 92 225
49 insufficienza respiratoria respiratory insufficiency respiratory insufficiency 724 1060
50 venosa; insufficienza venous; insufficiency venous insufficiency 30324 71150
51 venosa; arti; inferiori; insufficienza; sospetta venous; limbs; lower; insufficiency; suspected venous insufficiency 19532 28831
52 superficiale; venosa; insufficienza superficial; venous; insufficiency venous insufficiency 276 613
53 insufficienza; venosa; cronica venous; insufficiency; chronic venous insufficiency 415 506
54 venosa; safenectomia; insufficienza venous; saphenectomy; insufficiency venous insufficiency 280 311
55 venosa; claudicatio; insufficienza venous; claudication; insufficiency venous insufficiency 109 180
56 venosa; arti; insufficienza; varici venous; limb; insufficiency; varicose venous insufficiency 110 112
57 insufficienza; venosa; arti; inferiori venous; insufficiency; limbs; lower venous insufficiency 23 102
58 venoso; circolo; deficit; arti; venous; circulation; insufficiency; limbs venous insufficiency 74 102
59 venosa; insufficienza; arti; inferiori; edemi venous; insufficiency; limbs; lower; edemas venous insufficiency 47 58
60 insufficienza; venosa venous; insufficiency venous insufficiency 0 0
61 arti; venosa; inferiori limbs; venous; lower venous insufficiency 0 0
62 venosa; edemi; declivi; insufficienza venous; edemas; declivity; insufficiency venous insufficiency 0 0
63 venosa; insufficienza; edemi; arti; inferiori venous; insufficiency; edemas; limbs; lower venous insufficiency 0 0
64 venosa; cronica venous; chronic venous insufficiency 0 0
65 insufficienza; venosa; cronica; sospetta venous; insufficiency; chronic; suspected venous insufficiency 264 374
66 intervento; ginocchio; artroprotesi; valutazione surgery; knee; arthroplasty; evaluation intervention 67 77
67 lista; trapianto; rene; inserimento list; transplant; kidney; insertion intervention 61 69
68 lista; trapianto; valutazione; inserimento; rene; sinistra; dialisi; intervento list; transplant; evaluation; insertion; kidney; left; dialysis; intervention intervention 49 52
69 ipercolesterolemia hypercholesterolaemia hypercholesterolaemia 363 416
70 ipoestesia hypoesthesia hypoesthesia 31 80
71 ischemia ischaemia ischaemia 517 617
72 parestesie paresthesias paresthesias 1834 3965
73 parestesie; gamba; sinistra paresthesias; leg; left paresthesias 541 891
74 pesantezza heaviness heaviness 215 475
75 arti; inferiori; linfedema; pesantezza limbs; lower; lymphedema; heaviness heaviness 25 124
76 arti; inferiori; pesantezza limbs; lower; heaviness heaviness 18 106
77 safenectomia saphenectomy saphenectomy 1988 4177
78 safenectomia; esiti saphenectomy; outcomes saphenectomy 1374 3617
79 safenectomia; pregressa saphenectomy; previous saphenectomy 342 379
80 safenectomia; edemi saphenectomy; edemas saphenectomy 41 117
81 sindrome metabolica metabolic syndrome metabolic syndrome 12 48
82 stent stent stent 301 612
83 tromboflebite thrombophlebitis thrombophlebitis 7073 17415
84 flebite; sospetta phlebitis; suspected thrombophlebitis 5408 12379
85 flebite; gamba; superficiale phlebitis; leg; superficial thrombophlebitis 3037 8850
86 varicoflebite; sospetta; tromboflebite varicophlebitis; suspected; thrombophlebitis thrombophlebitis 4546 7739
87 tromboflebite; superficiale thrombophlebitis; superficial thrombophlebitis 950 1648
88 gamba; tromboflebite; destra leg; thrombophlebitis; right thrombophlebitis 397 486
89 tromboflebite; safena thrombophlebitis; saphenous vein thrombophlebitis 386 384
90 flebite; coscia; superficiale phlebitis; thigh; superficial thrombophlebitis 109 194
91 flebopatia; sospetta; arti; inferiori; sospetta phlebopathy; suspected; lower; limbs thrombophlebitis 209 175
92 tromboflebite; poplitea; arti; inferiori thrombophlebitis; popliteal; lower; limbs thrombophlebitis 54 102
93 flebite; superficiale; gamba phlebitis; superficial; leg thrombophlebitis 48 97
94 tromboflebite; sospetta; thrombophlebitis; suspected thrombophlebitis 30 68
95 trombosi thrombosis thrombosis 17060 43558
96 trombosi; venosa; profonda thrombosis; venous; deep thrombosis 17453 38230
97 trombosi; vascolare; thrombosis; vascular thrombosis 5086 10180
98 trombosi; venosa venous; thrombosis thrombosis 1325 3337
99 venosa; trombosi; profonda; sospetto; segni; clinici; clinico venous; thrombosis; deep; suspected; signs; clinical thrombosis 566 1152
100 poplitea; trombosi popliteal; thrombosis thrombosis 607 642
101 venosa; trombosi; poplitea venous; thrombosis; popliteal thrombosis 407 525
102 trombosi; safena; insufficienza thrombosis; saphenous; insufficiency thrombosis 362 472
103 segni; clinici; sospetto; clinico; trombosi; venosa; profonda; edema signs; clinical; suspected; clinical; venous; thrombosis; deep; edema thrombosis 342 389
104 safena; trombosi saphenous vein; thrombosis thrombosis 301 263
105 post; trombotica post; thrombotic thrombosis 36 73
106 tumefazione swelling swelling 533 741
107 tumefazione; gamba swelling; leg swelling 172 448
108 destro; polpaccio; tumefazione; paziente right; calf; swelling; patient swelling 35 49
109 ulcera; vascolare ulcer; vascular ulcer 828 3861
110 ulcera ulcer ulcer 1258 2411
111 varici varices varices 13999 35939
112 arti; inferiori; varici limbs; lower; varices varices 7018 18157
113 venose; varici venous; varices varices 4957 13554
114 varicosita; stasi varicosities; stasis varices 957 2701
115 varici; arti; inferiori varices; limbs; lower limbs varices 219 277
116 vene; varicose; algie veins; varicose; algie varices 121 194
117 safena; varici saphenous veins; varicose varices 47 136
118 varicosita; varicosita’ varicosities; varicositis varices 86 97
119 vascolari; esiti; polmonite; varici; chirurgici; linfangite vascular; outcomes; pneumonia; varices; surgical; lymphangitis varices 60 88
120 varici; sospette varicose veins; suspected varices 34 63
121 varicosita; gamba varicosities; leg varices 80 63
122 varici; intervento varicosities; surgery varices intervention 3144 7126
123 vasculopatia; periferica vasculopathy; peripheral vasculopathy 3820 8200
124 vasculopatia vasculopathy; peripheral vasculopathy 769 1886
125 vasculopatia; diabetico vasculopathy; diabetic vasculopathy 393 408
126 arteriopatia; obliterante arteriopathy; obliterative vasculopathy 288 272
127 arteriopatia; obliterante; periferica arteriopathy; obliterative; peripheral vasculopathy 47 117
128 iperteso; vasculopatico; vasculopatia; dislipidemico hypertensive; vasculopathy; dyslipidemic vasculopathy 17 97
129 aneurisma; aorta; esiti; popliteo; addominale; aortico aneurysm; aorta; popliteal; abdominal; aortic vasculopathy 42 78
130 vasculopatia; arti; inferiori vasculopathy; limbs; lower vasculopathy 36 78
131 vasculopatia; arti; inferiori; polidistrettuale vasculopathy; limb; lower; polydistrict vasculopathy 35 49
132 vasculopatia; venosa vasculopathy; venous vasculopathy 25 34
133 arteriopatia arteriopathy vasculopathy 20 25
134 natura; determinare; periferica nature; determine; peripheral other 3978 10039
135 natura; determinare; arti; inferiori nature; determine; limbs; lower other 5018 9601
136 paziente; polsi; periferici patient; wrists; peripheral other 2018 8015
137 arti; inferiori limbs; lower other 2190 5165
138 safena; incontinenza saphenous; incontinence other 288 782
139 ipertrofia hypertrophy hypertrophy 506 591
140 tromboembolia; polmonare; esiti thromboembolism; pulmonary; outcomes other 386 379
141 embolia; polmonare embolism; pulmonary embolism 110 174
142 femoro; femorale; paziente; popliteo; sospetta; asse; esiti; venosa femoral; femoral; patient; popliteal; suspected; axis; outcomes; venous other 49 49
143 stenosi; venoso stenosis; venous other 30 49
144 dislipidemia; ipertensione dyslipidaemia; hypertension dyslipidaemia 25 44
145 aorta; sottorenale aorta; subrenal other 25 29
Table C.2: *
Table C.3: Summary of all clusters identified for FEC, including their keyword representation, the label assigned to them and the number of distinct clinical questions (CQ) and referrals (REF) that they contain.
ID CLUSTER KEYWORDS [ITA] CLUSTER KEYWORDS [ENG] CLUSTER LABEL # CQ # REF
1 addominalgia; alvo abdominalgia; alvo abdominal pain 163 371
2 addominalgie; natura; determinare abdominalgia; nature; determine abdominal pain 714 5.801
3 addominali; algie; dolori abdominalgia; algia; pain abdominal pain 894 5.430
4 addominali; coliche abdominal; colic abdominal pain 240 3.775
5 alvo; addominalgie bowel; abdominal pain abdominal pain 224 2.507
6 calo; ponderale; addominalgia fall; weight; abdominalgia abdominal pain 100 215
7 natura; determinare; coliche nature; determine; colic abdominal pain 325 2.590
8 natura; determinare; addominali; coliche nature; determine; abdominal; colic abdominal pain 3.687 5.605
9 fianco; dolore; persistente flank; pain; persistent abdominal pain 61 121
10 fossa; addominalgia fossa; abdominalgia abdominal pain 192 2.125
11 fossa; iliaca; algia fossa; iliac; algia abdominal pain 90 194
12 fossa; iliaca; algie fossa; iliac; algia abdominal pain 33 71
13 fossa; iliaca; dolore fossa; iliac; pain abdominal pain 302 662
14 ipocondrio; destra; dolore; sinistra hypochondrium; right; pain; left abdominal pain 127 281
15 quadranti; addominalgia; addominalgie quadrant; abdominalgia; abdominalgia abdominal pain 33 74
16 stipsi; addominalgia constipation; abdominalgia abdominal pain 501 1.105
17 stipsi; addominali constipation; abdominal abdominal pain 112 250
18 stipsi; addominali; algie constipation; abdominal; pain abdominal pain 61 139
19 alvo; alterazione; alterazioni alvo; alteration altered bowel 432 3.500
20 alvo; alterno bowel; abdominalgia altered bowel 610 2.496
21 alvo; alterno; addominalgia bowel; abdominalgia altered bowel 287 1.925
22 alvo; calo; ponderale bowel; decline; weight altered bowel 196 1.655
23 alvo; stipsi bowel; constipation altered bowel 129 277
24 fossa; alvo fossa; alvo altered bowel 197 1.080
25 anemia; emoglobina anaemia; haemoglobin anaemia 61 128
26 anemia; natura; determinare; grave anaemia; nature; determine; severe anaemia 602 4.392
27 anemia; sangue; occulto; positivo anaemia; blood; occult; positive anaemia 315 2.366
28 anemia; sideropenica; origine; indeterminata anaemia; sideropenic; origin; undetermined anaemia 332 6.663
29 anemia; sideropenica; sangue; occulto anaemia; sideropenic; blood; occult anaemia 49 112
30 anemizzazione; grave; severa; sangue anaemia; severe; blood anaemia 456 3.814
31 anemizzazione; natura; determinare anaemia; nature; determine anaemia 250 2.504
32 sideropenica; anemia iron deficiency; anaemia anaemia 365 3.154
33 diarrea; colite; cronica diarrhoea; colitis; chronic chronic diarrhoea 95 757
34 diarrea; cronica diarrhoea; chronic chronic diarrhoea 778 7.668
35 diarroico; alvo diarrhoeic; alvo chronic diarrhoea 193 2.694
36 natura; determinare; diarrea; dolore; addominale nature; determine; diarrhoea; pain; abdominal chronic diarrhoea 409 2.414
37 colite; ischemica colitis; ischaemic chronic inflammatory bowel disease 119 286
38 crohn; morbo crohn disease; disease chronic inflammatory bowel disease 540 3.812
39 crohn; morbo; ileo crohn; disease; ileum chronic inflammatory bowel disease 71 202
40 infiammatoria; malattia; intestinale; cronica inflammatory; disease; intestinal; chronic chronic inflammatory bowel disease 220 2.333
41 malattie; infiammatorie; croniche; intestinali diseases; inflammatory; chronic; intestinal; disease chronic inflammatory bowel disease 313 3.780
42 malattie; infiammatorie; croniche; intestinali; calprotectina inflammatory; disease; chronic; intestinal; calprotectin chronic inflammatory bowel disease 19 51
43 stipsi; colite constipation; colitis chronic inflammatory bowel disease 74 185
44 ulcerosa; colite ulcerative; colitis chronic inflammatory bowel disease 2.985 7.323
45 ulcerosa; rettocolite ulcerative; rectocolitis chronic inflammatory bowel disease 283 3.498
46 adenocarcinoma; colon adenocarcinoma; colon colon cancer 336 753
47 adenoma; adenomi; colon adenoma; adenomas; colon colon cancer 50 102
48 alvo; tumore; colon bowel; cancer; colon colon cancer 57 120
49 anamnesi; colon anamnesis; colon colon cancer 76 301
50 colico; tumore colic; cancer colon cancer 97 219
51 colon; adenoma; adenomi; linfoma colon; adenoma; lymphoma colon cancer 123 394
52 colon; discendente; tumore colon; descending; cancer colon cancer 61 137
53 colon; pregressa; asportazione; tumore colon; previous; removal; cancer colon cancer 41 80
54 colon; ricerca; resezione; dopo; sincrone; cancro; lesioni; curativa colon; research; resection; after; synchronous; cancer; lesions; curative colon cancer 171 2.878
55 ematochezia; tumore; colon haematochezia; cancer; colon colon cancer 389 867
56 emicolectomia; tumore; esiti; destra; colon haemolectomy; cancer; outcomes; right; colon colon cancer 452 1.055
57 neoformazione; rettale; retto; colon neoformation; rectal; rectum; colon colon cancer 51 119
58 neoplasia; colon neoplasia; colon colon cancer 438 1.323
59 neoplasia; eteroplasia neoplasm; heteroplasm colon cancer 113 289
60 sigma; tumore sigma; cancer colon cancer 197 449
61 stipsi; tumore; colon constipation; cancer; colon colon cancer 286 682
62 tumore; colon cancer; colon colon cancer 5.725 16.757
63 tumore; colon; controllo cancer; colon; control colon cancer 108 324
64 tumore; colon; famigliarita cancer; colon; familial colon cancer 346 1.111
65 tumore; intestinale cancer; intestine colon cancer 305 840
66 tumore; pregresso; colon cancer; proctorragia; colon colon cancer 306 815
67 tumore; proctorragia; colon cancer; proctorragia; colon colon cancer 71 148
68 tumore; retto cancer; rectum colon cancer 266 671
69 ostinata; stipsi obstinate; constipation constipation 2.356 5.962
70 stipsi; cronica constipation; chronic constipation 745 7.521
71 stipsi; fossa; iliaca constipation; iliac fossa constipation 27 59
72 stipsi; ostinata; addominali constipation; obstinate; abdominal constipation 88 190
73 stipsi; ponderale; calo constipation; ponderal; fall constipation 122 1.377
74 diverticolite diverticulitis diverticulitis 759 3.148
75 diverticolite; acuta; recente diverticulitis; acute; recent diverticulitis 373 1.172
76 sigma; diverticolite sigma; diverticulitis diverticulitis 101 221
77 sigmoidite; sospetta; recente; episodio; diverticolare; esiti; diverticolite; diverticolosi sigmoiditis; suspected; recent; episode; diverticular; outcome; diverticulitis; diverticulosis diverticulitis 55 128
78 alvo; diverticolosi bowel; diverticulosis diverticulosis 359 2.126
79 diverticolare; malattia diverticular; disease diverticulosis 147 438
80 diverticoli; diverticolare; colon; poliposi diverticula; diverticular; colon; polyposis diverticulosis 388 1.849
81 diverticolosi; addominalgia diverticulosis; abdominalgia diverticulosis 355 782
82 diverticolosi; addominali diverticulosis; abdominal diverticulosis 264 2.213
83 diverticolosi; colon diverticulosis; colon diverticulosis 337 3.281
84 diverticolosi; colon; diverticolite diverticulosis; colon; diverticulitis diverticulosis 253 754
85 diverticolosi; sigma diverticulosis; sigma diverticulosis 117 256
86 pregressa; asportazione; diverticolosi previous; removal; diverticulosis diverticulosis 412 2.096
87 pregressa; polipectomia; diverticolosi previous; polypectomy; diverticulosis diverticulosis 142 306
88 stipsi; diverticolosi constipation; diverticulosis diverticulosis 136 957
89 emorroidaria; sindrome; patologia haemorrhoidal; syndrome; pathology haemorrhoidal disease 113 259
90 prevenzione; emorroidi cancer prevention / familiarity; haemorrhoids haemorrhoidal disease 962 4.647
91 prolasso; emorroidario prolapse; haemorrhoids haemorrhoidal disease 182 473
92 lynch; sindrome; colon lynch; syndrome; colon lynch syndrome 140 787
93 addominalgia; proctorragia abdominalgia; proctorragia bleeding 99 217
94 alvo; ematochezia bowel; haematochezia bleeding 116 251
95 alvo; sangue; occulto; fecale bowel; blood; occult; faecal bleeding 175 382
96 alvo; sanguinamento; irregolare; digestivo bowel; bleeding; irregular; digestive bleeding 703 4.915
97 ematiche; perdite bleeding; leakage bleeding 151 639
98 ematochezia; positivo; sanguinamento; natura; determinare haematochezia; positive; bleeding; nature; determine bleeding 760 12.945
99 ematochezia; addominalgia; addominalgie haematochezia; abdominalgia; abdominalgia bleeding 61 134
100 ematochezia; ematochezie haematochezia; haematochezia bleeding 500 5.307
101 melena; riferita; episodio melaena; reported; episode bleeding 182 953
102 nuovi; anticoagulanti; orali; terapia; proctorragia; anemia; paziente new; anticoagulants; oral; therapy; proctorragia; anaemia; patient bleeding 66 152
103 occulto; alvo; sangue; positivo occult; alvo; blood; positive bleeding 41 86
104 occulto; fecale; sangue; anemia occult; faecal; blood; anaemia bleeding 824 1.917
105 occulto; sangue occult; blood bleeding 4.459 12.755
106 occulto; sangue; positivo occult; blood; positive bleeding 33 75
107 oncologica; prevenzione oncological; cancer prevention / familiarity bleeding 2.346 4.752
108 proctorragia proctorrhagia bleeding 279 1.236
109 proctorragia; diverticolosi proctorrhagia; diverticulosis bleeding 152 328
110 proctorragia; emorroidi; episodi proctorrhagia; haemorrhoids; episodes bleeding 571 1.439
111 proctorragia; rettorragia proctorragia; rectorrhagia bleeding 4.219 13.288
112 proctorragie; emorroidi proctorragia; haemorrhoids bleeding 65 220
113 prolasso; proctorragia; emorroidario prolapse; proctorragia; haemorrhoid bleeding 46 98
114 rettorragia; colon rectorrhagia; colon bleeding 48 104
115 rettorragia; episodi rectorrhagia; episodes bleeding 408 1.167
116 rettorragia; tumore; colon rectorrhagia; cancer; colon bleeding 74 157
117 sangue; episodi; presenza blood; episodes; presence bleeding 330 1.647
118 sangue; occulto blood; occult bleeding 263 619
119 sangue; occulto; anemizzazione; positivo blood; occult; anaemia; positive bleeding 122 292
120 sangue; occulto; calprotectina; fecale; positivo blood; occult; calprotectin; faecal; positive bleeding 25 64
121 sangue; occulto; fecale blood; occult; faecal bleeding 2.525 7.258
122 sangue; occulto; fecale; anemizzazione blood; occult; faecal; anaemia bleeding 224 494
123 sangue; occulto; fecale; calo; ponderale blood; fecal; occult; weight-loss bleeding 140 312
124 sangue; occulto; fecale; positivita blood; fecal; occult; positive bleeding 198 529
125 sangue; occulto; fecale; positivo; terapia; anemia blood; faecal; occult; positive; therapy; anaemia bleeding 38 96
126 sangue; occulto; fecale; tumore blood; occult; faecal; cancer bleeding 171 364
127 sangue; occulto; polipectomia blood; occult; polypectomy bleeding 105 250
128 sangue; occulto; positivo blood; occult; positive bleeding 69 151
129 sangue; stipsi; occulto; fecale blood; constipation; occult; faecal bleeding 163 357
130 stipsi; ematochezia constipation; haematochezia bleeding 91 190
131 stipsi; proctorragia constipation; proctorragia bleeding 156 347
132 adenomatosi; polipi adenomatous; polyps polyps 49 116
133 asportato; colon; polipo removed; colon; polyp polyps 182 424
134 asportazione; polipo removal; polyp polyps 457 1.052
135 colon; polipi colon; polyps polyps 2.079 6.733
136 exeresi; polipi; pregressa exeresis; polyps; previous polyps 71 174
137 follow; up; polipectomia follow; up; polypectomy polyps 133 333
138 polipectomia; colon polypectomy; colon polyps 89 236
139 polipectomia; controllo polypectomy; control polyps 516 2.263
140 polipectomia; endoscopica; pregressa polypectomy; endoscopic; previous polyps 109 356
141 polipectomia; pregressa; addominalgia polypectomy; previous; abdominalgia polyps 49 111
142 polipectomia; sigma polypectomy; sigma polyps 80 177
143 polipectomia; tumore; colon polypectomy; cancer; colon polyps 32 65
144 polipi; adenomi; sorveglianza; pazienti; serrati polyps; adenomas; surveillance; patients; serratus polyps 932 23.390
145 polipi; colon polyps; colon polyps 610 15.287
146 polipi; controllo polyps; control polyps 365 4.950
147 polipi; intestinali polyps; intestinal polyps 186 881
148 polipi; pregressi polyps; previous polyps 333 1.128
149 polipo; adenomatoso polyps; adenomatous polyps 44 104
150 polipo; anale polyp; anal polyps 123 340
151 polipo; colon polyp; colon polyps 168 5.634
152 poliposi; colon polyposis; colon polyps 711 1.847
153 poliposi; intestinale polyposis; intestinal polyps 231 1.876
154 poliposi; pregressa polyposis; previous polyps 53 119
155 poliposi; sospetta polyposis; suspected polyps 263 875
156 pregressa; polipectomia previous; polypectomy polyps 981 6.133
157 pregressa; poliposi previous; polyposis polyps 412 2.208
158 pregressa; sigma; asportazione; polipo previous; sigma; removal; polyp polyps 79 165
159 pregresso; polipo previous; polyp polyps 322 3.204
160 proctorragia; poliposi; colon proctorragia; polyposis; colon polyps 98 228
161 sigma; polipo sigma; polyp polyps 272 621
162 stipsi; ostinata; poliposi constipation; obstinate; polyposis polyps 38 81
163 alvo; familiarita; tumore alvo; familial; cancer cancer prevention / familiarity 212 604
164 anamnesi; tumore; colon; familiare anamnesis; cancer; colon; familial cancer prevention / familiarity 200 3.358
165 colon; familiarita; tumore colon; familial; cancer cancer prevention / familiarity 302 6.530
166 diagnosi; precoce; prestaz; diagnost; tumori early; diagnosis; performance; diagnostics; cancers cancer prevention / familiarity 106 1.378
167 familiarita; tumore; colon familial; cancer; colon cancer prevention / familiarity 1.137 3.286
168 famigliarita; tumore; colon familial; cancer; colon cancer prevention / familiarity 151 388
169 familiarita; colon; neoplasia familiarity; colon; neoplasm cancer prevention / familiarity 1.160 3.450
170 familiarita; prevenzione familiarity; cancer prevention / familiarity cancer prevention / familiarity 306 1.114
171 oncologica; familiarita; prevenzione; tumore oncological; familial; cancer prevention / familiarity; cancer cancer prevention / familiarity 195 453
172 prestaz; diagnost; diagnosi; precoce; tumori prior; diagnosis; early; diagnosis; cancers cancer prevention / familiarity 55 136
173 prestaz; diagnost; diagnosi; precoce; tumori; tumore; familiarita; colon service; diagnosis; early; diagnosis; cancers; cancer; familial; colon cancer prevention / familiarity 192 418
174 prestaz; diagnost; diagnosi; sangue; occulto; precoce; tumori performance; diagnosis; diagnosis; occult; blood; early; cancers cancer prevention / familiarity 49 126
175 tumore; anamnesi; familiare cancer; history; family history cancer prevention / familiarity 280 3.603
176 tumore; colon; familiarita cancer; colon; familial cancer prevention / familiarity 618 4.379
177 tumore; colon; familiarita’ cancer; colon; familial cancer prevention / familiarity 401 5.010
178 tumore; colon; proctorragia; familiarita cancer; colon; proctorship; familiarity cancer prevention / familiarity 127 274
179 tumore; familiarita cancer; familiarity cancer prevention / familiarity 3.709 8.696
180 tumore; intestino; familiarita cancer; intestine; familiarity cancer prevention / familiarity 64 412
181 asportazione; adenoma; esiti removal; adenoma; outcome surgery 45 91
182 esiti; resezione; retto outcomes; resection; rectum surgery 57 47
183 pregressa; asportazione previous; removal surgery 921 105
184 pregressa; asportazione; adenoma previous; removal; adenoma surgery 180 293
185 pregressa; emicolectomia previous; hemicolectomy surgery 75 85
186 renale; trapianto; lista; insufficienza; cronica; dialisi renal; transplantation; list; failure; chronic; dialysis surgery 25 64
187 trapianto; fegato transplant; liver surgery 86 112
188 colica; sospetta; lesione; riscontrata; altra; tecnica; indagine colic; suspected; lesion; found; other; technique; investigation suspected colic lesion 179 3.671
189 calo; ponderale fall; weight other 333 1.022
190 colon; irritabile colon; irritable other 263 1.432
191 displasia; adenoma; tubulare; grado dysplasia; adenoma; tubular; grade other 148 316
192 fossa; iliaca fossa; iliac other 24 54
193 grado; displasia grade; dysplasia other 84 191
194 prolasso; rettale prolapse; rectal other 57 122
195 sigma; ispessimento sigma; thickening other 43 87
196 subocclusione; intestinale sub-occlusion; intestinal other 81 202